### v2

# When does bottom-up beat top-down in hierarchical community detection?

Abstract
Hierarchical clustering of networks consists in finding a tree of communities in order that higher depths of the hierarchy reveal finer-grained community structures. There are two main classes of algorithms for this problem. Divisive (top-down) algorithms recursively partition the nodes into two (or more) communities, until a stopping rule indicates that no further split is needed. In contrast, agglomerative (bottom-up) algorithms first identify the smallest community structure and then repeatedly merge the communities by using a linkage method. In this article, we establish theoretical guarantees for recovering the hierarchy in the community structure of a hierarchical stochastic block model (HSBM) by a bottom-up algorithm. We also establish that this bottom-up algorithm attains the information-theoretic threshold for exact recovery at intermediate depths of the hierarchy. Notably, these recovery conditions are less restrictive compared to those existing for top-down algorithms. This shows that bottom-up algorithms extend the feasible region for achieving exact recovery at intermediate depths. Numerical experiments on both synthetic and real data sets confirm the superiority of bottom-up algorithms over top-down algorithms. In particular, a notable drawback of top-down algorithms is their tendency to produce dendrograms with inversions. These findings contribute to a better understanding of hierarchical clustering techniques and their applications in network analysis.
Keywords: Community detection, hierarchical clustering, linkage, stochastic block model.
1 Introduction
A system of pairwise interactions among entities can conveniently be represented by a graph, where the system entities are the nodes and the interactions are the edges. Data collected in such form is increasingly abundant in many disciplines, such as sociology, physics, economics, and biology New (18). Finding community structures by grouping nodes with similar connection patterns into clusters is one of the most important statistical analysis tasks on networks For (10); AD (22). Community structures are often hierarchical. For example, in a co-authorship network, we can partition the researchers, based on their primary discipline (such as mathematics, physics, computer science, etc); each of these fields can be further split into specific sub-disciplines. The sub-division of larger communities into smaller ones provides a finer division of the network.
Hierarchical communities can be inferred using top-down approaches, where the process begins by identifying the largest communities at the top of the hierarchy. These communities are then recursively decomposed into smaller sub-communities until a stopping rule indicates no further division is necessary. To identify the communities at the top of the hierarchy, one strategy is to progressively remove edges with the highest edge-betweenness centrality GN (02) or with the lowest edge-clustering coefficient RCC+ (04). Another strategy is to bi-partition the network by using spectral clustering  DHKM (06); BXKS (11); LLB+ (22). Top-down algorithms, however, possess several limitations. First, any clustering errors initially made become locked in and propagate to the next rounds of subdivisions, potentially compromising the accuracy of the predicted hierarchy. Second, the recursive splittings can overlook valuable information by ignoring the edges between communities that have already been separated. As a result, these approaches might not fully capture the interconnections between communities across different depths of the hierarchy.
In contrast, agglomerative (or bottom-up) algorithms take a different approach by constructing the hierarchy from the bottom upwards. These algorithms recursively merge smaller communities to form larger ones. Some bottom-up algorithms, such as the one proposed in BCGH (18); PL (05); CHCL (11), generate a complete dendrogram. A dendrogram is a tree whose leaves are individual nodes, whose branches and internal nodes represent merged clusters, and the lengths of the branches are proportional to the degree of similarity between the merged clusters. Other algorithms, such as greedy schemes maximizing modularity BGLL (08); HSH+ (10), produce partial dendrograms because the leaves are no longer associated with individual nodes but with the bottom communities. Alternatively, initially inferring the bottom flat communities using a graph-clustering algorithm and then merging these communities, through a procedure known as linkage CAKMTM (19), also produces a partial dendrogram. This last approach offers flexibility in building the dendrogram, by selectively incorporating community merges based on specific criteria.
Establishing theoretical guarantees by assessing the performance of HCD algorithms on random graph models is important to compare them. The hierarchical stochastic block model (HSBM) is a general model of random graphs with hierarchical communities. This model defines the hierarchical community structure as a rooted binary tree whose leaves correspond to primitive communities. Each node belongs to a primitive community, and the interactions between two nodes belonging to communities $a$ and $b$ depend only on the lowest common ancestor between $a$ and $b$ on the hierarchical tree.
We first establish that recovering the hierarchy of an HSBM is possible using a bottom-up algorithm when the average degree of the graph is only $\omega(1)$. Earlier studies required stronger degree growth conditions, such as $\Theta(N)$ in BXKS (11), $\omega(N^{1/2}\log^{p}N)$ (with $p=1/2$ in CAKMT (17) and $p=2$ in LTA+ (16)), or $\omega(\log^{1+p}N)$ (with $p=5$ for DHKM (06) and $p=1+\Omega(1)$ for LLB+ (22)).
In the context of hierarchical communities, it is possible to study community recovery at different depths of the hierarchy. We rigorously establish the information-theoretic threshold for the exact recovery of the communities at any intermediate depth of the hierarchy. Exact recovery is the strongest notion of recovery, as it is defined by the ability to fully recover the communities when the number of nodes $N$ tends to infinity. In particular, we establish that exact recovery at a higher depth is more challenging than exact recovery at a lower depth.
We also demonstrate that a bottom-up algorithm using the linkage procedure achieves this threshold. Previous research in this area showed that recursive spectral bi-partitioning, a top-down algorithm, recovers the depths 00 to $\ell$ of the hierarchy, asymptotically exactly if a set of conditions is respected LLB+ (22). Our work reveals that these conditions are stricter than the information-theoretic threshold. As a result, bottom-up algorithms extend the feasible region of exact recovery at intermediate depths, up to the information-theoretic limits.
Finally, in our numerical experiments, we employ a bottom-up algorithm that follows a two-step process. To infer the underlying primitive communities, we first apply a spectral algorithm on the graph’s Bethe-Hessian matrix  SKZ (14); DCT (21). Next, we build the hierarchy using the linkage procedure. To provide a comprehensive comparison, we evaluate this bottom-up algorithm against recursive spectral bi-partitioning Lei (20); LLL (21), as it is the most relevant top-down algorithm whose performance has been theoretically established and numerically validated. Our findings on synthetic data sets demonstrate that the bottom-up algorithm achieves exact recovery at intermediate depths, up to the information-theoretic thresholds, whereas recursive spectral bi-partitioning fails to do so. Furthermore, we show that the dendrogram produced by recursive spectral bi-partitioning can suffer from inversions. Inversions occur when the clustering algorithm incorrectly places a lower-depth cluster above a higher-depth cluster in the dendrogram, thus distorting the true hierarchical structure. Such inversions can lead to misleading interpretations of the hierarchical relationships within the data.
The paper is structured as follows. In Section 2, we describe top-down and bottom-up approaches for hierarchical community detection. The HSBM is defined in Section 3, where we also derive conditions for recovering the hierarchy by the linkage procedure, as well as its robustness to mistakes made in the initial clustering. Next, we investigate the exact recovery at intermediate depths of the hierarchy in Section 4. We discuss these results in light of the existing literature in Section 5. Finally, Section 6 is devoted to the numerical experiments.
Notations
We denote by $[N]$ the set $\{1,\cdots,N\}$, by $\operatorname{Ber}(p)$ a Bernoulli random variable with parameter $p$, and by $\mathrm{Multi}(N,\mathcal{S},\pi)$ a multinomial distribution with $N$ trials, a (finite, discrete) set of possible outcomes $\mathcal{S}$, and a vector of probability $\pi$ such that $\pi_{s}$ is the probability of success of $s\in\mathcal{S}$. The Frobenius norm of a matrix is denoted $\|\cdot\|_{F}$.
The Rényi divergence of order $t\in(0,1)$ of a Bernoulli distribution $\operatorname{Ber}(p)$ from another distribution $\operatorname{Ber}(q)$ is defined as $\operatorname{D}_{t}\left(\operatorname{Ber}(p)\|\operatorname{Ber}(q)\right)=\frac{1}{t-1}\log\left((1-p)^{t}(1-q)^{1-t}+p^{t}q^{1-t}\right)$. For $t=1/2$, we write $\operatorname{D}_{1/2}\left(\operatorname{Ber}(p),\operatorname{Ber}(q)\right)=-2\log\left(\sqrt{(1-p)(1-q)}+\sqrt{pq}\right)$ and this quantity is symmetric in $p$ and $q$.
We focus on undirected graphs $G=(V,E)$ whose node set is $V=[N]$ and adjacency matrix $A=\left(A_{ij}\right)\in\{0,1\}^{N\times N}$. For a subset $V_{1}$ of the node set $V$, we let $G[V_{1}]$ be the subgraph of $G$ induced by $V_{1}$. We denote by $\mathcal{N}_{\mathcal{T}}$ (resp., $\mathcal{L}_{\mathcal{T}}$) the internal nodes (resp., the leaves) of a tree $\mathcal{T}$. For an internal node $u$ of a rooted $\mathcal{T}$, denote by $\mathcal{T}[u]$ the sub-tree rooted at $u$. We denote $\mathrm{lca}(u,v)$ the lowest common ancestor between two nodes $u,v\in\mathcal{T}$.
2 Hierarchical Community Detection
Many networks present a hierarchical community structure. The primitive communities at the bottom of the hierarchy are a collection of subsets $\mathcal{C}=(C_{1},\cdots,C_{K})$ that partition the original node set $V$ into $K$ disjoint sets. These primitive communities are the leaves of a rooted tree $\mathcal{T}$, and this tree defines the hierarchical relationship between the communities. This section reviews two main strategies for hierarchical community detection (HCD).
2.1 Divisive (top-down) Algorithms
Divisive (top-down) HCD algorithms begin with one single community containing all the nodes. This community is recursively split until a selection rule indicates that no further splits are needed. This can be summarized as follows:
  1. 1.
    apply a selection rule to decide if the community contains sub-communities. If no, stop; if yes, split into two communities;
  2. 2.
    recursively repeat step 1 on each of the two sub-communities found.
Different choices for the stopping rule or the bi-partitioning algorithms have been explored DHKM (06); BXKS (11); LLB+ (22).
Each recursive splitting of the graph loses some information. Indeed, consider the two clusters $C_{0}$ and $C_{1}$ obtained after the first split. The next step splits $C_{0}$ (resp., $C_{1}$) into two clusters $C_{00}$ and $C_{01}$ (resp., $C_{10}$ and $C_{11}$), based only on the induced sub-graph $G[C_{0}]$ (resp., $G[C_{1}]$). This implies that the clustering of $C_{0}$ does not take into account the edges from $C_{0}$ to $C_{1}$. If the edge densities between $C_{00}$ and $C_{1}$ are different from the edge densities between $C_{01}$ and $C_{1}$, then this valuable information is not used by a top-down algorithm.
Furthermore, it is important to note that the resulting tree is unweighted. Although it is possible to compute similarities between pairs of predicted clusters (for example using the edge density between two clusters), these similarities are not guaranteed to increase with the depth of the dendrogram; inversions can occur. These inversions, as we will emphasize in the numerical section, lead to discrepancies in the hierarchical structure portrayed by the dendrogram.
2.2 Agglomerative (bottom-up) Algorithms
Agglomerative (bottom-up) HCD algorithms construct a sequence of clusterings in an ascending manner, where the dendrogram is created from its leaves to its root. The initial clustering corresponds to the leaves of the dendrogram, and the hierarchy is progressively built up by iteratively merging the most similar clusters.
A first approach is to initially assign each node to a separate community and then merge the clusters that minimize a distance metric New (04); PL (05); CHCL (11); BCGH (18). Although these methods generate a complete dendrogram, determining the depth in the hierarchy where the community structure becomes meaningful is an old yet challenging problem Moj (77); GBW (24).
A second approach directly estimates the bottom clusters by using a flat (non-hierarchical) graph-clustering algorithm. These bottom clusters are then sequentially merged two by two, based on a similarity measure. This can be summarized as follows:
  1. 1.
    apply a graph-clustering algorithm to find the bottom clusters;
  2. 2.
    1. (a)
      compute the similarity between all pairs of bottom clusters;
    2. (b)
      merge the two clusters that are the most similar [^1] When measuring the similarity between pairs of clusters by using edge density, it is implicitly assumed that the clustering structure exhibits assortativity. This means that nodes within the same community are more likely to be connected. However, in cases where the communities are disassortative, the focus shifts towards merging clusters that display the lowest similarity, thus reflecting the lack of intra-community connections., and update the similarities between this new cluster and the existing ones;
    3. (c)
      repeat step (2b) until all clusters have been merged into a single one.
The abundant literature on graph clustering provides numerous candidate algorithms for the first stage. We note that, in most practical applications, the true number of bottom clusters $K$ is often unknown and needs to be inferred at this stage as well. In this paper, we use the edge density as the similarity measure between pairs of clusters. The edge density between two node sets $V_{1},V_{2}\subset V$ is defined by
$\rho\left(V_{1},V_{2}\right)\ =\ \frac{w\left(V_{1},V_{2}\right)}{|V_{1}|\cdot|V_{2}|}\quad\text{ where }\quad w\left(V_{1},V_{2}\right)\ =\ \sum_{i\in V_{1}}\sum_{j\in V_{2}}A_{ij},$ (2.1)
Recomputing the similarities in Step 2b is done as follows. Let $\widehat{C}_{1},\cdots,\widehat{C}_{\widehat{K}}$ be the clusters initially predicted by the flat graph-clustering algorithm. Suppose that at some step of the algorithm, clusters $\widehat{C}_{k_{1}}$ and $\widehat{C}_{k_{2}}$ are the most similar and hence are merged to give a new cluster $\widehat{C}_{k_{1}\cup k_{2}}=\widehat{C}_{k_{1}}\cup\widehat{C}_{k_{2}}$. We have for $\ell\not\in\{k_{1},k_{2}\}$,
$\rho\left(\widehat{C}_{k_{1}\cup k_{2}},\widehat{C}_{\ell}\right)\ =\ \frac{\left|\widehat{C}_{k_{1}}\right|}{\left|\widehat{C}_{k_{1}\cup k_{2}}\right|}\rho\left(\widehat{C}_{k_{1}},\widehat{C}_{\ell}\right)+\frac{\left|\widehat{C}_{k_{2}}\right|}{\left|\widehat{C}_{k_{1}\cup k_{2}}\right|}\rho\left(\widehat{C}_{k_{2}},\widehat{C}_{\ell}\right).$ (2.2)
Therefore, the edge density naturally defines an average-linkage procedure for merging the clusters identified through the initial flat clustering. In particular, this procedure guarantees that the resulting dendrogram will be free from inversions, thus ensuring a coherent hierarchical structure MC (17). Algorithm 1 provides a concise summary of this process.
Input: Graph $G=(V,E)$, flat graph-clustering algorithm flat-clustering.
Output: Hierarchical clustering tree $\widehat{\mathcal{T}}$ and its associated bottom clusters $\widehat{C}_{1},\cdots,\widehat{C}_{\widehat{K}}$.
1 Let $\widehat{K}$ be the number of clusters and $\widehat{\mathcal{C}}=\left(\widehat{C}_{1},\cdots,\widehat{C}_{\widehat{K}}\right)$ the clusters inferred by flat-clustering; For all $k,\ell\in\left[\widehat{K}\right]$, compute $\rho\left(\widehat{C}_{k},\widehat{C}_{\ell}\right)$ as in (2.1);
while $\left|\widehat{\mathcal{C}}\right|\geq 2$ do:
  • Let $\widehat{C}_{k},\widehat{C}_{\ell}\in\operatorname*{arg\,max}\limits_{C\neq C^{\prime}\in\widehat{\mathcal{C}}}\rho\left(C,C^{\prime}\right)$;
  • Let $\widehat{\mathcal{C}}\leftarrow\widehat{\mathcal{C}}\,\backslash\,\left\{\widehat{C}_{k},\widehat{C}_{\ell}\right\}$;
  • For any $C\in\widehat{\mathcal{C}}$ compute $\rho\left(\widehat{C}_{k\cup\ell},C\right)$ as defined in (2.2);
  • Let $\widehat{\mathcal{C}}\leftarrow\widehat{\mathcal{C}}\cup\left\{\widehat{C}_{k\cup\ell}\right\}$.
Return: The bottom communities $\widehat{\mathcal{C}}$ and the tree $\widehat{\mathcal{T}}$ capturing the sequence of nested merges in the while-loop.
Algorithm 1 Bottom-up hierarchical community detection.
3 Tree Recovery from the Bottom
In this section, we study the asymptotic performance of bottom-up algorithms on a class of random graphs with hierarchical community structures. We define the model in Section 3.1 and, in Section 3.2, we establish the condition for recovering the hierarchical tree from the bottom communities. In Section 3.3, we show that the average-linkage procedure is robust to errors made in the initial clustering step.
3.1 Hierarchical Stochastic Block Model
The hierarchical stochastic block model (HSBM) is a class of random graphs whose nodes are partitioned into latent hierarchical communities. Before defining this model formally, let us introduce some notations. Each node $u$ of a rooted binary tree $\mathcal{T}$ is represented by a binary string as follows. The root is indexed by the empty string $\emptyset$. Each non-root node $u$ of the tree is labeled by the unique binary string $u=u_{1}u_{2}\cdots u_{q}\cdots$ that records the path from the root (i.e., $u_{q}=1$ if step $q$ of the path is along the right branch of the split and $u_{q}=0$ otherwise). The depth of node $u$ is denoted by $|u|$ and coincides with its distance from the root [^2] By convention, $|\emptyset|=0$.. Finally, with this parametrization, the lowest common ancestor $\mathrm{lca}(u,v)$ of two nodes $u,v\in\mathcal{T}$ is simply the longest common prefix of the binary strings $u$ and $v$.
Here, we denote by $\mathcal{L}_{\mathcal{T}}$ the leaves of the tree $\mathcal{T}$. We will assign each node of the graph $G$ to one leaf of $\mathcal{T}$ and denote by $C_{a}$ the set of nodes assigned to the leaf $a$. This forms the primitive communities $\mathcal{C}=\left(C_{a}\right)_{a\in\mathcal{L}_{\mathcal{T}}}$. Any internal node $u$ of the tree is associated with a super-community $C_{u}$ such that $C_{u}=\cup_{a\in\mathcal{L}_{\mathcal{T}[u]}}C_{a}$ where $\mathcal{L}_{\mathcal{T}[u]}$ denotes the leaves of the sub-tree of $\mathcal{T}$ rooted at $u$. In particular, we have $C_{\emptyset}=V$ and $C_{v}\subset C_{u}$ if $v$ is a child of $u$.
We suppose that the probability $p_{ab}$ of having a link between two nodes belonging to the primitive communities $C_{a}$ and $C_{b}$ depends only on the lowest common ancestor of $a$ and $b$. Hence, we denote the probability $p_{ab}$ by $p(\mathrm{lca}(a,b))$. In an assortative setting, we have $p(u)<p(v)$ if $v$ is a child of $u$. Finally, we need to rule out settings in which the difference $p(v)-p(u)$ decreases too fast with respect to $p(u)$. To avoid this flatness of $p$ issue, we always assume that lim $\liminf\frac{\left|p(v)-p(u)\right|}{\max\{p(u),p(v)\}}>0$ for any child $v$ of a parent $u$. We now define an assortative HSBM.
Definition 1.
Let $N$ be a positive integer, $\mathcal{T}$ a rooted binary tree with $|\mathcal{L}_{\mathcal{T}}|=K$ leaves, $\pi=(\pi_{a})_{a\in\mathcal{L}_{\mathcal{T}}}$ a probability vector and a function $p\colon\mathcal{T}\to[0,1]$ such that for $u,v\in\mathcal{T}$ we have $p(u)<p(v)$ (assortativity) if $v$ is a child of $u$. A hierarchical stochastic block model (HSBM) is a graph $G=(V,E)$ such that $V=[N]$ and
  1. 1.
    each node $i\in[N]$ is independently assigned to a community $C_{a}$ where $a$ is sampled from $[K]$ according to $\pi$;
  2. 2.
    two nodes $i\in C_{a}$ and $j\in C_{b}$ are connected with probability $p(\mathrm{lca}(a,b))$.
An important particular case of HSBM is the binary tree SBM (BTSBM), in which the tree $\mathcal{T}$ is full and balanced, and the probability of a link between two nodes in clusters $C_{a}$ and $C_{b}$ depends only on the depth of $\mathrm{lca}(a,b)$, i.e., $p(\mathrm{lca}(a,b))=p_{|\mathrm{lca}(a,b)|}$ for all $a,b\in\mathcal{L}_{\mathcal{T}}$. In particular, the number of communities of a BTSBM is $K=2^{d_{\mathcal{T}}}$, and assortativity implies that $p_{0}<p_{1}<\cdots<p_{d_{\mathcal{T}}}$. We illustrate an HSBM and a BTSBM in Figure 1.
(a) Hierarchical stochastic block model.
(b) Binary tree SBM of depth 3.
Figure 1: Examples of an HSBM (a) and a BTSBM (b), with the binary string representation of each node. The link probabilities are $p(u)$ for the HSBM and $p_{|u|}$ for the BTSBM. The grey-colored rectangles represent the super-communities.
3.2 Tree Recovery from the Bottom Clusters
In this section, we study the recovery of the hierarchical tree $\mathcal{T}$ and the bottom communities $\mathcal{C}=\left(C_{\ell_{1}},\cdots,C_{\ell_{K}}\right)$ associated with the $K$ leaves of $\mathcal{T}$. For an estimator $\widehat{\mathcal{C}}=\left(\widehat{C}_{1},\cdots,\widehat{C}_{K}\right)$ of $\mathcal{C}$ verifying $|\mathcal{C}|=|\widehat{\mathcal{C}}|=K$, the number of mis-clustered nodes is
$\mathrm{loss}\left(\mathcal{C},\widehat{\mathcal{C}}\right)\ =\ \min_{\tau\in\mathcal{S}_{[K]}}\sum_{k=1}^{K}\left|C_{\ell_{k}}\,\Delta\,\widehat{C}_{\tau(k)}\right|,$ (3.1)
where $B\,\Delta\,C=\left(B\cup C\right)\backslash\left(B\cap C\right)$ denotes the symmetric difference between two sets $B,C\subset V$. The minimum is taken over the set $\mathcal{S}_{[K]}$ of all permutations of $[K]$, because we recover only the bottom communities up to a global permutation of the community labels. We study sequences of networks indexed by the number of nodes $N$ and for which the interaction probabilities might depend on $N$. An estimator $\widehat{\mathcal{C}}$ of $\mathcal{C}$ achieves almost exact recovery if $\mathbb{P}\left(\mathrm{loss}\left(\mathcal{C},\widehat{\mathcal{C}}\right)=o(N)\right)=1-o(1)$ and exact recovery if $\mathbb{P}\left(\mathrm{loss}\left(\mathcal{C},\widehat{\mathcal{C}}\right)=0\right)=1-o(1)$.
The number of edges between two communities $C_{a}$ and $C_{b}$ is binomially distributed with mean $|C_{a}|\cdot|C_{b}|\cdot p_{ab}$. Therefore, if $p_{ab}=\omega(N^{-2})$, this binomial random variable is concentrated around its mean and we have $\rho(C_{a},C_{b})=(1+o(1))p_{ab}$. This suggests that, once the primitive communities are almost exactly recovered, the average-linkage procedure successfully recovers the tree from its leaves if the edge probability between different communities is $\omega\left(N^{-2}\right)$. The proof is more involved because (i) we compute $\rho(\widehat{C}_{a},\widehat{C}_{b})$ and not $\rho(C_{a},C_{b})$ and (ii) the estimator $\widehat{\mathcal{C}}$ is correlated with the graph structure. In particular, to establish that $\rho(\widehat{C}_{a},\widehat{C}_{b})=(1+o(1))p_{ab}$, we impose $p_{ab}=\omega(N^{-1})$.
Theorem 1.
Consider an assortative HSBM whose latent binary tree $\mathcal{T}$ has $K$ leaves. Suppose that $K$ is constant, $\min_{a\in\mathcal{L}_{\mathcal{T}}}\pi_{a}>0$, and $p(t)=a(t)\delta_{N}$ where $(a(t))_{t\in\mathcal{T}}$ are positive constants (independent of $N$) and $N\delta_{N}=\omega(1)$. Let $\widehat{\mathcal{C}}$ be an estimator of $\mathcal{C}$, possibly correlated with the graph edges and such that $|\widehat{\mathcal{C}}|=|\mathcal{C}|$. If $\widehat{\mathcal{C}}$ is almost exact, then the average-linkage procedure correctly recovers $\mathcal{T}$ (starting from $\widehat{\mathcal{C}}$).
We prove Theorem 1 in Appendix A.1. The assumption of the existence of an almost exact estimator $\widehat{\mathcal{C}}$ of $\mathcal{C}$ is not limiting here. Indeed, under the assumptions of Theorem 1, we can obtain such an estimator by the spectral algorithm of YP (16) or by the agnostic-degree-profiling algorithm of AS15b .
3.3 Robustness of the Linkage Procedure
In this section, we study the robustness of the linkage procedure to classification errors made by the initial clustering, when the number of errors scale as $\Theta(N)$. For simplicity, we restrict the study to BTSBMs.
Suppose that each node has a probability $\eta$ to be misclustered and that misclusterings occur uniformly across all cluster pairs. More precisely, the probability that a node $i\in[N]$ belonging to block $a$ is misclustered to a different block $b\neq a$ depends only on the depth of the lowest common ancestor between $a$ and $b$. We denote this probability by $\zeta(|\mathrm{lca}(a,b)|)$. Finally, let $\widetilde{\mathcal{C}}(\mathcal{C},\zeta)$ be the bottom clusters obtained by this procedure [^3] Note that $\widetilde{\mathcal{C}}(\mathcal{C},\zeta)$ depends on $\zeta$ and $\mathcal{C}$ but is independent of the graph edges. This is unrealistic because the linkage procedure starts from the predicted bottom clusters found by applying a clustering algorithm on the graph (see Algorithm 1). From a theoretical perspective, we can split the original graph $G=(V,E)$ into two subgraphs $G_{1}=(V,E_{1})$ and $G_{2}=(V,E_{2})$ such that $E_{1}\cup E_{2}=E$. An edge $(u,v)\in E$ is in $E_{1}$ with probability $\gamma$, otherwise it is in $E_{2}$. We apply the flat clustering algorithms on $G_{1}$ to obtain $\widehat{\mathcal{C}}^{1}$ and run the linkage procedure on $G_{2}$; $\widehat{\mathcal{C}}^{1}$ is independent of $E_{2}$..
A natural assumption is that nodes are more likely to be misclustered in blocks close to their true block; this is verified if $\zeta$ is monotonically non-decreasing.
Theorem 2.
Let $G$ be a BTSBM whose latent binary tree $\mathcal{T}$ has depth $d$ and $K=2^{d}$ leaves. Suppose that $\pi_{a}=1/K$ for all $a\in\mathcal{L}_{\mathcal{T}}$ and that $\min\limits_{u\in\mathcal{T}}p(u)=\omega(N^{-2})$. Assume that $\zeta$ is non-decreasing and verifies $(\zeta(d-1)+\zeta(d-2))/2<1-\eta$. Let $\widetilde{\mathcal{C}}=\widetilde{\mathcal{C}}(\mathcal{C},\zeta)$ be a clustering obtained as described at the beginning of Section 3.3. Then the average-linkage procedure correctly recovers $\mathcal{T}$ (starting from $\widetilde{\mathcal{C}}$).
Example 1 shows that if (i) the misclustered nodes are assigned to a cluster uniformly at random among all $K$ clusters and if (ii) the proportion $1-\eta$ of correctly clustered nodes is strictly larger than $1/K$, then average-linkage recovers the tree. As $1/K$ is the expected proportion of correctly clustered nodes resulting from a random guess, Condition (ii) is weak.
Example 1.
Let $\zeta_{\mathrm{uniform}}(h)=\eta/(K-1)$ for $10\leq h\leq d-$ 1. If $1-\eta>1/K$, then the average-linkage recovers $\mathcal{T}$ starting from $\widetilde{\mathcal{C}}_{\mathrm{uniform}}=\widetilde{\mathcal{C}}(\mathcal{C},\zeta_{\mathrm{uniform}})$.
In an adversarial setting, $\zeta$ might not be non-decreasing. For instance, consider a scenario where misclustered nodes are deliberately assigned to a cluster chosen uniformly at random from those farthest away from their true cluster. Define $\widetilde{\mathcal{C}}_{\mathrm{adversarial}}=\widetilde{\mathcal{C}}(\mathcal{C},\zeta_{\mathrm{adversarial}})$ where
$\zeta_{\mathrm{adversarial}}(h)\ =\ \begin{cases}\frac{\eta}{2^{d-1}}&\text{ if }h=0,\\ 0&\text{ if }h\in\{1,\cdots,d-1\}.\end{cases}$ (3.2)
Under additional conditions given in Proposition 1, the average-linkage procedure recovers the tree from $\widetilde{\mathcal{C}}_{\mathrm{adversarial}}$.
Proposition 1.
Let $G$ be a BTSBM whose latent binary tree $\mathcal{T}$ has depth $K=\Theta(1)$ leaves. Suppose that $\pi_{a}=1/K$ for all $a\in\mathcal{L}_{\mathcal{T}}$ and that $\min\limits_{u\in\mathcal{T}}p(u)=\omega(N^{-2})$. Suppose that $\eta<1/2$. Then the average-linkage procedure correctly recovers $\mathcal{T}$ (starting from $\widetilde{\mathcal{C}}_{\mathrm{adversarial}}$) if one of the following conditions is verified:
$(i)\ \bar{p}_{1}<p_{d-1}\qquad\text{ or }\qquad(ii)\ \bar{p}_{1}\geq p_{d-1}\text{ and }\eta<\eta_{-}.$
where $\bar{p}_{1}=\frac{1}{2^{d-1}}\left(p_{d}+\sum_{k=1}^{d-1}2^{k-1}p_{d-k}\right)$ and $\eta_{-}=\frac{p_{d-1}+\bar{p}_{1}-2p_{0}-\sqrt{(\bar{p}_{1}-p_{d-1})(\bar{p}_{1}-p_{0})}}{p_{d-1}+3\bar{p}_{1}-4p_{0}}$.
4 Exact Recovery at Intermediate Depths
We now focus on the exact recovery of communities at intermediate depths within the hierarchy and establish tight information-theoretic conditions determining the feasibility of achieving exact recovery at a specific depth. It is intuitively expected that recovering super-communities of higher depths should be comparatively easier than the recovery of lower depths. Therefore, we expect scenarios where the exact recovery of the primitive communities might be unattainable, but where the exact recovery of super-communities at intermediate depths remains achievable. To provide context, we initially recapitulate key findings on exact recovery in non-hierarchical stochastic block models (SBMs) in Section 4.1. We then present our main results in Section 4.2, which gives the precise conditions for exact recovery at intermediate depths.
4.1 Chernoff-Hellinger Divergence
The hardness of separating nodes that belong to a primitive community $a\in\mathcal{L}_{\mathcal{T}}$ from nodes in community $b\in\mathcal{L}_{\mathcal{T}}$ is quantified by the Chernoff-Hellinger divergence, denoted by $\mathrm{CH}(a,b)=\mathrm{CH}(a,b,\pi,p,\mathcal{T})$ and defined by444 AS15a originally defined the Chernoff-Hellinger divergence as $\mathrm{CH}_{AS}(a,b)\ =\ \sup_{t\in(0,1)}(1-t)\sum_{c\in\mathcal{L}_{\mathcal{T}}}\pi_{c}\left(tp_{ac}+(1-t)p_{bc}-p_{ac}^{t}p_{bc}^{1-t}\right).$ When $p_{ab}=o(1)$, we note that $(1-t)\operatorname{D}_{t}\left(\operatorname{Ber}\left(p_{ac}\right)\|\operatorname{Ber}\left(p_{bc}\right)\right)\ =\ (1+o(1))\left(tp_{ac}+(1-t)p_{bc}-p_{ac}^{t}p_{bc}^{1-t}\right)$, hence our definition of $\mathrm{CH}(a,b)$ in (4.1) coincides with the original definition $\mathrm{CH}_{AS}(a,b)$ (up to second-order terms).
$\mathrm{CH}(a,b)\ =\ \sup_{t\in(0,1)}(1-t)\sum_{c\in\mathcal{L}_{\mathcal{T}}}\pi_{c}\operatorname{D}_{t}\left(\operatorname{Ber}\left(p_{ac}\right)\|\operatorname{Ber}\left(p_{bc}\right)\right),$ (4.1)
where $\operatorname{D}_{t}$ denotes the Rényi divergence of order $t$. The key quantity assessing the possibility or impossibility of exact recovery in an HSBM is the minimal Chernoff-Hellinger divergence between all pairs of clusters. We denote it by $I=I(\pi,p,\mathcal{T})$, and it is defined by
$I\ =\ \min_{\begin{subarray}{c}a,b\in[K]\\ a\neq b\end{subarray}}\ \mathrm{CH}(a,b).$ (4.2)
The condition for achieving exact recovery of the communities displays a phase transition. More precisely, when $p_{ab}=\Theta(\log N/N)$, exact recovery of the primitive communities is possible if $\lim\frac{NI}{\log N}>1$ and is impossible when $\lim\frac{NI}{\log N}<1$ AS15a ; AS15b .
Example 2.
Consider a BTSBM with $K=2^{d}$ communities and $\pi_{a}=1/K$ for all $a\in[K]$. Suppose that for all $t\in[d]\colon p_{t}=a_{t}\log N/N$, with $a_{t}$ positive constants. Simple computations (see for example (Lei, 20, Lemma 6.6)) yield that
$I\ =\ \frac{1}{K}\operatorname{D}_{1/2}\left(\operatorname{Ber}(p_{d}),\operatorname{Ber}(p_{d-1})\right)\ =\ \frac{1+o(1)}{K}\left(\sqrt{a_{d}}-\sqrt{a_{d-1}}\right)^{2}\,\frac{\log N}{N}.$
Exact recovery of $\mathcal{C}$ is possible if $\left(\sqrt{a_{d}}-\sqrt{a_{d-1}}\right)^{2}>K$. This condition involves only $a_{d}$ and $a_{d-1}$, but not $a_{d-2},a_{d-3},\cdots,a_{0}$.
4.2 Exact recovery at Intermediate Depths
Although the exact recovery of the primitive communities is feasible when the quantity $I$, as defined in Equation (4.2), exceeds the threshold of $\log N/N$, this insight alone does not provide the requirements for achieving exact recovery of the super-communities of higher depth within the hierarchy. In this section, we answer this question by determining the precise information-theoretic threshold for achieving the exact recovery of super-communities at a specific depth. Furthermore, we demonstrate that a bottom-up algorithm successfully attains this threshold.
We say that a node $u$ of $\mathcal{T}$ is at depth $q$ if its distance from the root is $q$, or if $u$ is a leaf [^5] The second condition ($u$ is a leaf) is only needed if the binary tree is not perfect, as we highlight in Example 3.. We denote by $\mathcal{S}_{q}$ (resp., $\mathcal{S}_{\leq q}$) the set of nodes at depth $q$ (resp., at a depth less than or equal to $q$), i.e.,
$\mathcal{S}_{q}\ =\ \Big{\{}u\in\mathcal{T}\colon\big{(}|u|=q\big{)}\text{ or }\big{(}|u|\leq q\text{ and }u\text{ is a leaf}\big{)}\Big{\}}\quad\text{ and }\quad\mathcal{S}_{\leq q}\ =\ \bigcup\limits_{r\leq q}\mathcal{S}_{r}.$
The set of super-communities at depth $q$ is hence
$\mathrm{sc}\left(q,\mathcal{C},\mathcal{T}\right)\ =\ \left\{C_{u}\right\}_{u\in\mathcal{S}_{q}}\quad\text{ with }\quad C_{u}\ =\ \bigcup_{a\in\mathcal{L}_{\mathcal{T}}\colon a_{1:q}=u}C_{a}.$ (4.3)
Example 3.
For the tree of Figure 1(a) we have
$\mathrm{sc}\left(q,\mathcal{C},\mathcal{T}\right)\ =\ \begin{cases}\{C_{0},C_{1}\}&\text{ if }q=1,\\ \{C_{00},C_{01},C_{10},C_{11}\}&\text{ if }q=2,\\ \{C_{00},C_{010},C_{011},C_{10},C_{11}\}&\text{ if }q=3.\end{cases}$ (4.4)
Because the tree of Figure 1(a) is not full and complete, the community $C_{10}$ is present at depth $2$ and at depth $3$. In contrast, for the tree of Figure 1(b), we have
$\mathrm{sc}\left(q,\mathcal{C},\mathcal{T}\right)\ =\ \begin{cases}\{C_{0},C_{1}\}&\text{ if }q=1,\\ \{C_{00},C_{01},C_{10},C_{11}\}&\text{ if }q=2,\\ \{C_{000},C_{001},C_{010},C_{011},C_{100},C_{101},C_{110},C_{111}\}&\text{ if }q=3.\end{cases}$
The recovery of super-communities at depth $q$ can be affected by mistakes made at depth $q^{\prime}>q$. For example, if we consider the super-communities given in (4.4), then the recovery at depth $2$ of $C_{01}$ can be exact, even if nodes belonging to $C_{010}$ are mistakenly classified in $C_{011}$. But, this is not the case if nodes in $C_{010}$ are mistakenly classified in $C_{00}$. As a result, the exact recovery of the super-communities at depth $2$ might still be achievable even if the communities at depth $3$ cannot be recovered, as long as the errors occur within a super-community. This can occur, for instance, if $C_{010}$ and $C_{011}$ are very hard to separate, whereas $C_{010}$, $C_{00}$, $C_{10}$ and $C_{11}$ are easy to separate. We recall from Section 3.2 that the hardness to distinguish two communities is expressed in terms of a Chernoff-Hellinger divergence (4.1). The difficulty in separating the primitive communities that do not belong to the same super-community at depth $q$ is quantified by the minimum Chernoff-Hellinger divergence taken across all pairs of primitive communities that do not belong to the same super-community at depth $q$. This is the quantity $I_{q}=I(q,\pi,p,\mathcal{T})$ defined as
$I_{q}\ =\ \min_{\begin{subarray}{c}a\neq b\in\mathcal{L}_{\mathcal{T}}\\ \mathrm{lca}(a,b)\in\mathcal{S}_{\leq q-1}\end{subarray}}\mathrm{CH}\left(a,b\right),$ (4.5)
where the condition $\mathrm{lca}(a,b)\in\mathcal{S}_{\leq q-1}$ ensures that the lowest common ancestor of $a$ and $b$ has a depth less than or equal to $q-1$. When $q=d$, the minimum in Equation (4.5) is taken over all pairs of primitive communities, and the divergence $I_{d}$ is equal to the divergence $I$ defined in (4.2). Theorem 3 states the condition for recovering the communities at depth $q$.
Theorem 3.
Let $G$ be an HSBM whose latent binary tree $\mathcal{T}$ has $K=2^{d}$ leaves. Let $q\in\{1,\cdots,d_{\mathcal{T}}\}$ and denote by $I_{q}$ the quantity defined in (4.5). The following holds:
  1. (i)
    exact recovery of the super-communities at depth $q$ is impossible if lim $\limsup\frac{NI_{q}}{\log N}<1$;
  2. (ii)
    if lim $\liminf\frac{NI_{q}}{\log N}>1$ and $\min\limits_{u\in\mathcal{T}}p(u)=\omega(N^{-2})$, then the agnostic-degree-profiling algorithm of AS15b exactly recovers the super-communities at depth $q$.
Because $I_{q}$ is the minimum taken over all $\mathcal{S}_{\leq q-1}$, the quantity $I_{q}$ is non-decreasing in $q$. This is the reason why recovery at a higher intermediate depth $q$ is easier than recovery at a lower depth $q^{\prime}>q$. An extra condition $\min\limits_{u\in\mathcal{T}}p(u)=\omega(N^{-2})$ is present in point (ii) of Theorem 3 to rule out some particular cases [^6] For example, consider a toy BTSBM with 3 layers for which $p_{0}=0$, $p_{1}=N^{-10}$, $p_{2}=2\log N/N$ and $p_{3}=2.01\log N/N$. Although exact recovery at depth $3$ is impossible, exact recovery at depth $2$ is still possible. Nonetheless, exact recovery at depth 1 is impossible because, with a high probability, there are no edges between the super-communities having depth $2$..
We prove Theorem 3 in Appendix B.1. Although the quantity $I_{q}$ defined in (4.5) is generally hard to simplify, the following lemma provides a simple expression for BTSBMs.
Lemma 1.
For a BTSBM with $K=2^{d}$ balanced communities ($\pi_{a}=1/K$ for all $a\in\mathcal{L}_{\mathcal{T}}$), the minimum Chernoff-Hellinger divergence defined in Equation (4.5) is
$I_{q}\ =\ \frac{1}{K}\left(\operatorname{D}_{1/2}(\operatorname{Ber}(p_{q-1}),\operatorname{Ber}(p_{d}))+\sum_{k=1}^{d-q}2^{k-1}\operatorname{D}_{1/2}\left(\operatorname{Ber}(p_{q-1}),\operatorname{Ber}(p_{d-k})\right)\right).$
We prove Lemma 1 in Appendix B.2. In particular, Lemma 1 shows that when $q=d$, we have $I_{d}=K^{-1}\operatorname{D}_{1/2}\left(\operatorname{Ber}(p_{d-1}),\operatorname{Ber}(p_{d})\right)$, and hence we recover the divergence $I$ defined in (4.2) (see also Example 2).
5 Discussion
5.1 Previous Work on Exact Recovery in HSBM
Early works by Dasgupta et al. DHKM (06) establish conditions for the exact recovery by top-down algorithms in an HSBM. Nonetheless, the recovery is ensured only in relatively dense regimes (specifically, average degree at least $\log^{6}N$), and the algorithm requires an unspecified choice of hyper-parameters. Balakrishnan et al. BXKS (11) show that a simple recursive spectral bi-partitioning algorithm recovers the hierarchical communities in a class of hierarchically structured weighted networks, where the weighted interactions are perturbed by sub-Gaussian noise. Nonetheless, the conditions require again a dense regime ((BXKS, 11, Assumption 1) states that all weights are strictly positive, hence the average degree scales as $\Theta(N)$), and the proposed algorithm has no stopping criterion. Recent work by Li et al. LLB+ (22) demonstrates that the same recursive spectral bi-partitioning algorithm exactly recovers the communities, when the average degree grows as $\Omega(\log^{2+\epsilon}N)$ for some $\epsilon>0$ (this condition can be further relaxed to $\Omega(\log N)$ by a refined analysis Lei (20)). The analysis of LLB+ (22) also allows for an unbounded number of communities and provides a consistent stopping criterion.
The linkage++ algorithm proposed by Cohen-Addad et al. CAKMT (17); CAKMTM (19) is a bottom-up algorithm that first estimates $K$ primitive communities by SVD and then successively merges them by using a linkage procedure. As SVD requires the number of clusters $K$ as an input, the authors propose to run the algorithms for every $K=2,\cdots,O(\log N)$ and to choose the optimal $K$ that leads to the hierarchical tree with the smallest Dasgupta-cost Das (16). Moreover, their analysis requires a dense setting in which the average degree grows as $\Omega(\sqrt{N\log N})$.
5.2 Top-Down HCD and Exact Recovery at Intermediate Depth
Sufficient conditions for the exact recovery of super-communities in BTSBMs by recursive spectral bi-partitioning algorithms were first established in LLB+ (22), under the assumptions that all connection probabilities have the same scaling factor $\delta_{N}$ (i.e., $p(t)=\Theta(\delta_{N})$ for all $t\in\mathcal{T}$) verifying $\delta_{N}=\Omega(\log^{2+\epsilon}N/N)$. The result was later refined in Lei (20) to remove the extra logarithmic factors and in LLL (21) to enable different scaling factors at different depths of the hierarchy. To the best of our knowledge, these are the only works involving recovery at intermediate depths of the hierarchy. To compare these results with Theorem 3, let us consider a BTSBM with $p(t)=a_{|t|}\log N/N$ where $a_{|t|}$ is constant independent of $N$. The average connection probability in a super-community of depth $q$ is then $\bar{a}_{q}\log N/N$ where $\bar{a}_{q}=\frac{1}{2^{d-q}}\left(a_{d}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}\right)$. Then (Lei, 20, Theorem 6.5) states that the recursive spectral bi-partitioning algorithm exactly recovers all the super-communities up to depth $\ell$ if
$\left(\sqrt{\bar{a}_{q}}-\sqrt{a_{q-1}}\right)^{2}>2^{q}\quad\text{ for all }q\in[\ell].$ (5.1)
The condition $\left(\sqrt{\bar{a}_{q}}-\sqrt{a_{q-1}}\right)^{2}>2^{q}$ corresponds to the exact recovery threshold of a symmetric SBM with $2^{q}$ communities and intra-community (resp., inter-community) link probabilities $\bar{a}_{q}\log N/N$ (resp., $a_{q-1}\log N/N$). Although the depth $q$ of the hierarchy is composed of $2^{q}$ super-communities with average intra-connection probability $\bar{a}_{q}\log N/N$ and inter-connection probability $a_{q-1}\log N/N$, Condition (5.1) does not match the information-theoretic threshold derived in Theorem 3. Indeed, the structure of a super-community at depth $q$ is not simply an Erdős-Rényi random graph with connection probability $\bar{a}_{q}\log N/N$, but is instead an SBM composed of $2^{d-q}$ primitive communities, each of them interacting differently with the rest of the network.
For all $q\in[d]$, let us define $J_{q}^{\mathrm{td}}$ and $J_{q}^{\mathrm{bu}}$ by
$J_{q}^{\mathrm{td}}$ $\ =\ \frac{1}{2^{d}}\left(\sqrt{a_{d}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}}-\sqrt{2^{d-q}a_{q-1}}\right)^{2},$ (5.2)
$J_{q}^{\mathrm{bu}}$ $\ =\ \frac{1}{2^{d}}\left(\left(\sqrt{a_{q-1}}-\sqrt{a_{d}}\right)^{2}+\sum_{k=1}^{d-q}2^{k-1}\left(\sqrt{a_{q-1}}-\sqrt{a_{d-k}}\right)^{2}\right).$ (5.3)
Condition (5.1) for exact recovery of the super-communities up to depth $\ell$ is equivalent to
$\min_{q\in[\ell]}J_{q}^{\mathrm{td}}>1,$ (5.4)
whereas the corresponding condition for a bottom-up algorithm given by Theorem 3(ii) and Lemma 1 can be recast as
$J_{\ell}^{\mathrm{bu}}>1.$ (5.5)
Let us compare (5.4) and (5.5). First, notice that for some choices of $a_{0},\cdots,a_{d}$, the function $q\mapsto J_{q}^{\mathrm{td}}$ given in (5.2) is not monotonically non-increasing [^7] Let $d=3$. For $(a_{0},a_{1},a_{2},a_{3})=(2.2,2.5,3,25)$ we have $(J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}})=(0.96,1.17,1.33)$, but for $(a_{0},a_{1},a_{2},a_{3})=(3,9,15,21)$ we have $(J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}})=(1.89,0.39,0.06)$. Finally, for $(a_{0},a_{1},a_{2},a_{3})=(2.2,2.4,4,22)$ we have $(J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}})=(0.85,1.02,0.90)$. Thus, the quantities $J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}}$ can be increasing, decreasing or interlacing., which contradicts the intuitive fact that the recovery at lower depths is harder than the recovery at higher depths. Moreover, although $J_{d}^{\mathrm{bu}}=J_{d}^{\mathrm{td}}$, Lemma 5 (deferred in Appendix B.3) shows that $J_{q}^{\mathrm{bu}}>J_{q}^{\mathrm{td}}$ for all $q\in[d-1]$, hence Condition (5.4) is strictly more restrictive than Condition (5.5). It is not clear whether the results of Lei (20) can be further improved as exact recovery by spectral bi-partitioning requires entry-wise concentration of the eigenvectors, which is established using sophisticated $\ell_{2\to\infty}$ perturbation bounds.
6 Numerical Results
In this section, we investigate the empirical performance of different HCD strategies on synthetic and real networks. More precisely, we compare the bottom-up approach of Algorithm 1 with other HCD algorithms. To implement Algorithm 1, we use spectral clustering with the Bethe-Hessian matrix, as proposed in DCT (21), as the initial clustering step. This is a spectral method that does not require the knowledge of the number of communities $K$ and that was shown to work very well on synthetic and real data sets [^8] The code for this algorithm is available online at https://lorenzodallamico.github.io/codes/.. In all the following, bottom-up refers to Algorithm 1, top-down to the recursive bi-partitioning algorithm of LLB+ (22). To facilitate reproducible research, our code is available on GitHub [^9] https://github.com/dakrod/bottom_up_HCD.
6.1 Synthetic Data Sets
We compare bottom-up and top-down approaches on synthetic data sets by computing the accuracy at different depths. We define accuracy at depth $q$ as $1-\mathrm{loss}\left(\mathrm{sc}(q,\mathcal{C},\mathcal{T}),\mathrm{sc}(q,\widehat{\mathcal{C}},\widehat{\mathcal{T}})\right)/N$, where $\mathrm{loss}$ is given by (3.1) and $\mathrm{sc}(q,\mathcal{C},\mathcal{T})$ (resp., $\mathrm{sc}(q,\widehat{\mathcal{C}},\widehat{\mathcal{T}})$) are the ground truth (resp., predicted) super-communities defined in (4.3).
6.1.1 Binary Tree SBMs
We first generate synthetic BTSBMs with depth $d$, where the probability of an edge between two nodes whose lowest common ancestor has depth $k$ is $p_{k}=a_{k}\log N/N$. We compare the accuracy of top-down and bottom-up at each depth. We show in Figure 2 the results obtained on BTSBMs with depth $3$, with $200$ nodes in each primitive community (thus $N=2^{3}\times 400=3200$ nodes in total). We let $a_{0}=40$, $a_{3}=100$, and the values of $a_{1}$ and $a_{2}$ vary in the range ($a_{0},a_{3}$), with the condition $a_{1}<a_{2}$. Solid lines in each panel show the exact recovery threshold of the given method at different depths. We observe the strong alignment between the theoretical guarantees and the results of the numerical simulations.
Figure 2: Performance of bottom-up and top-down algorithms on BTSBMs with depth 3, $N=3200$ nodes, and interaction probabilities $p_{k}=a_{k}\log N/N$ where $a_{0}=40$ and $a_{3}=100$, as a function of $a_{1}$ and $a_{2}$. We vary $a_{1}\leq a_{2}$ from $a_{0}$ to $a_{3}$. The empirical performance of the algorithms is measured by the accuracy at each depth, given by the color scale (results are averaged over 10 realizations). Large circles represent exact recovery (i.e., perfect accuracy on each of the 10 runs), and small crosses represent a non-exact recovery. The colored solid lines delimit the theoretical exact recovery thresholds for each algorithm on the various depths (given by Equations (5.4)-(5.5)); for a given depth $q$, these equations provide a single condition for bottom-up, but $q$ conditions for top-down. At depths 1 and 2, the regimes where exact recovery can be achieved are the areas above the solid line(s). At depth 3, the area lies below the threshold drawn by the red line (and above the blue and green lines for top-down; this area forms a small triangle).
6.1.2 Ternary Tree SBMs
As the hierarchical community structure cannot always be represented by a binary tree, we also perform experiments on ternary-tree stochastic block models with depth 3 (the ternary tree is drawn in Figure 3(a)), 100 nodes in each bottom cluster, and the probability of an edge between two nodes whose lowest common ancestor has depth $k$ is $p_{k}=a_{k}\log N/N$.
We first show in Figure 3 the dendrograms and trees obtained by top-down and bottom-up algorithms. Although both algorithms generate binary trees, ternary structures appear in Figures 3(b) and 3(c). Nevertheless, we observe in Figure 3(c) that the dendrograms obtained by the top-down algorithm show some inversions.
Next, we proceed as in Section 6.1.1, by fixing $a_{0}$ to 40 and $a_{3}$ to 100, and by varying the values of $a_{1}$ and $a_{2}$ from $a_{0}$ to $a_{3}$ (with the condition $a_{1}<a_{2}$). We observe in Figure 4 differences in the performances of top-down and bottom-up. Indeed, in this setting, the bottom-up algorithm recovers communities exactly up to the theoretical thresholds. Moreover, the accuracy obtained by top-down is lower than the one obtained by bottom-up.
(a) Ground truth tree
(b) bottom-up
(c) top-down
Figure 3: (a) A ternary tree of depth 3 used as ground truth. (b)-(c) Dendrograms obtained by bottom-up and top-down algorithms on a Ternary Tree SBM of depth 3, $N=2700$, and interaction probabilities $p_{k}=a_{k}\log N/N$ with $(a_{0},a_{1},a_{2},a_{3})=(10,30,40,130)$.
Figure 4: Performance of bottom-up and top-down algorithms on Ternary Tree SBMs of depth 3, $N=2700$ nodes, and interaction probabilities $p_{k}=a_{k}\log N/N$ with $a_{0}=10$ and $a_{3}=130$, as a function of $a_{1}$ and $a_{2}$. We vary $a_{1}\leq a_{2}$ from $a_{0}$ to $a_{3}$. The performance of the algorithms is measured by the accuracy at each depth (averaged over 10 realizations), and the exact recovery threshold at different depths is shown in coloured solid lines. Exact recovery is shown with large circles and non-exact recovery with small crosses.
6.1.3 Robustness of Linkage to Misclustering Errors
We numerically test the robustness of the average-linkage procedure, when the expected number of mistakes done in the clustering of the primitive communities is $\eta N$, with $\eta>0$. Consider the two following scenarios described in Section 3.3:
  • a uniform scenario in which $\widetilde{\mathcal{C}}_{\mathrm{uniform}}$ is sampled so that the misclustered nodes are assigned to a cluster uniformly at random among all clusters (this is the setting of Example 1);
  • an adversarial scenario in which $\widetilde{\mathcal{C}}_{\mathrm{adversarial}}$ is sampled so that the misclustered nodes are assigned to a cluster chosen uniformly at random among the clusters that are the furthest away from their true cluster (this is the setting of Proposition 1).
We denote by $\widehat{\mathcal{T}}_{\mathrm{uniform}}$ (resp., by $\widehat{\mathcal{T}}_{\mathrm{adversarial}}$) the tree obtained by the average-linkage algorithm from the bottom communities $\widetilde{\mathcal{C}}_{\mathrm{uniform}}$ (resp., from $\widetilde{\mathcal{C}}_{\mathrm{adversarial}}$). To assess the correct recovery of the tree, we denote by $S\left(\mathcal{T},\mathcal{C}\right)$ the $N$-by-$N$ tree similarity matrix. The elements of this matrix are defined, for $i\in C_{a}$ and $j\in C_{b}$ (with $a,b\in\mathcal{L}_{\mathcal{T}}$), by
$\Big{(}S\left(\mathcal{T},\mathcal{C}\right)\Big{)}_{ij}\ =\ |\mathrm{lca}_{\mathcal{T}}(a,b)|,$ (6.1)
where $|\mathrm{lca}_{\mathcal{T}}(a,b)|$ is the depth of the lowest common ancestor of $C_{a}$ and $C_{b}$ in the tree $\mathcal{T}$. The error in recovering the tree is then defined as
$r_{s}\ =\ \frac{\left\|S\left(\widehat{\mathcal{T}}_{s},\widetilde{\mathcal{C}}_{s}\right)-S\left(\mathcal{T},\widetilde{\mathcal{C}}_{s}\right)\right\|_{F}^{2}}{\left\|S\left(\mathcal{T},\widetilde{\mathcal{C}}_{s}\right)\right\|_{F}^{2}},$
where $s\in\{\text{uniform},\text{adversarial}\}$. This ratio equals 00, as long as the algorithm succeeds in restoring the tree completely, even if the bottom communities $\widetilde{\mathcal{C}}_{s}$ contain mistakes. We plot in Figure 5 the values of these ratios. As established in Example 1, we see that the tree is correctly recovered in the uniform scenario. For the adversarial scenario, we observe the failure of the tree recovery when the proportion of misclustered nodes $\eta$ becomes too large. Nonetheless, when $\eta$ is small, the average-linkage procedure still recovers the tree correctly. This highlights the robustness of the average-linkage and is in alignment with Proposition 1.
(a) Uniform mistakes
(b) Adversarial mistakes
Figure 5: Recovery from error done by the average-linkage procedure on a a balanced BTSBM with depth 3, $N=4000$, and $p_{k}=0.08\beta^{3-k}$, as a function of $\beta$ and of the misclustering error rate $\eta$. To distinguish exact recovery ($r_{s}=0$) from almost exact recovery, square symbols are used for the former, and circles for the latter. Results are averaged over 25 realizations.
6.2 Real Datasets
6.2.1 High-School Contact Dataset
We illustrate the performance of HCD algorithms on a data set of face-to-face contacts between high school students. The data set was collected from the high school Lycée Thiers in Marseilles, France MFB (15) and is available at http://www.sociopatterns.org/. The 9 communities are classes of 327 students, and the weighted interactions are the number of proximity encounters during 5 school days. We can expect this network to possess hierarchical relationships between communities. Indeed, although each student belongs to one of the nine classes, these classes also define the student’s specialty. More precisely, there are four specializations:“MP” stands for mathematics and physics, “PC” for physics and chemistry, “PSI” for engineering, and “BIO” for biology. Furthermore, there are three “MP” classes (MP, MP*1, and MP*2), two “PC” classes (PC and PC*), one “PSI” class (PSI*), and three “BIO” classes (2BIO1, 2BIO2, and 2BIO3). Therefore, we expect that the hierarchy reveals the students’ specialties at a higher depth and the students’ classes at a lower depth.
The results of the two HCD algorithms are shown in Figure 6. The bottom-up algorithm always detects $31$ communities. In contrast, out of 100 runs, the top-down algorithm predicts on average 8.93 communities (7 runs detect $8$ communities and the remaining 93 runs detect 9 communities). The predictions of both algorithms are aligned with the ground truth. Indeed, the adjusted mutual information (AMI) VEB (09) between the nine classes taken as ground truth labels and the predicted communities is $0.938$ for bottom-up [^10] We compute the AMI with respect to the configuration obtained after merging the 31 bottom communities into 9 super-communities. and $0.945$ for top-down. We observe that bottom-up successfully recovers the class and specialization structure but, in addition, finds smaller communities. This likely represents small groups of friends inside a class: the bottom-up algorithm reveals a community structure that is richer than the ground truth itself.
Figure 6: Bottom-up and top-down algorithms on the high school data set. Vertices correspond to the students, the colors to the true classes, and edges of the graph are in grey. The hierarchical tree is drawn in black, and its root is marked by a star symbol.
6.2.2 Power Grid
We next consider the power grid of continental Europe from the map of the Union for the Coordination of Transmission of Electricity (UCTE) [^11] http://www.ucte.org. We use the same data set that was previously used for HCD SDYB (12). Figures 7 and 8 show the outputs of the bottom-up and top-down, respectively. We observe a higher correlation to geographical positions in the output of bottom-up. Moreover, the dendrogram obtained by top-down shows a significant amount of inversions, contrary to the one obtained by bottom-up.
(a) 8 communities.
(b) 2 communities.
(c) Dendrogram.
Figure 7: Bottom-up algorithm on the power-grid network.
(a) 8 communities.
(b) 2 communities.
(c) Dendrogram.
Figure 8: Recursive spectral bi-partitioning algorithm on the power grid network.
7 Conclusion
Top-down approaches need to make partition decisions for large communities, without exploiting side information about the internal structure of these communities. In a sparse regime, finding a bi-partition is as difficult as finding a flat community structure. As a consequence, some nodes are misclassified, and these misclassifications are increasingly locked in as the algorithm progresses down to smaller communities. In contrast, a bottom-up approach inherently exploits lower-level structure, if present. At each step, a bottom-up algorithm needs to classify only the communities at the subsequent higher level, rather than classifying all the nodes individually. This is an easier problem (even if some errors are carried up), because (a) the number of classification decisions is much smaller, and (b) the number of edges available for each decision is much larger (number of edges between two lower-depth communities versus number of incident edges on an individual node).
In this paper, we have quantified this fundamental advantage within a class of random graph models (HSBM). We have proven that the latent tree of an HSBM can be recovered under weaker conditions than in the literature (average degree scaling as $\omega(1)$). Moreover, we have established that super-communities of intermediate depths could be exactly recovered up to the information-theoretic threshold, thus improving the previously known conditions for top-down algorithms. Finally, we have shown that the theoretic advantage of bottom-up carries over to relevant scales and real-world data. Both on synthetic and on real data sets, bottom-up HCD achieves better performance than top-down HCD.
Supplementary material
When does bottom-up beat top-down in hierarchical community detection?
Appendix A Proofs of Section 3
A.1 Proof of Theorem 1
Proof of Theorem 1.
With the assumption of this theorem, $\widehat{\mathcal{C}}$ is an almost exact estimator of the true clusters $\mathcal{C}$. Let $a\neq b\in\mathcal{L}_{\mathcal{T}}$, and recall the definition of the edge density $\rho(\cdot,\cdot)$ in (2.1). Lemma 2 implies that with high probability the edge density between the estimated clusters $\widehat{C}_{a}$ and $\widehat{C}_{b}$ is $\rho(\widehat{C}_{a},\widehat{C}_{b})=(1+o(1))p_{ab}$. Since the network is assortative and the hierarchy is non-flat, this implies that the first step of the linkage finds the two closest clusters, say $\widehat{C}_{a_{1}}$ and $\widehat{C}_{a_{2}}$. Then we see that for $b\in\mathcal{L}_{\mathcal{T}}\backslash\{a_{1},a_{2}\}$ we have $\mathrm{lca}(a_{1},b)=\mathrm{lca}(a_{2},b)$ and hence $p_{a_{1}b}=p_{a_{2}b}$. Therefore, from (2.2),
$\rho\left(\widehat{C}_{a_{1}\cup a_{2}},\widehat{C}_{b}\right)\ =\ \frac{\left|\widehat{C}_{a_{1}}\right|}{\left|\widehat{C}_{a_{1}}\right|+\left|\widehat{C}_{a_{2}}\right|}\rho\left(\widehat{C}_{a_{1}},\widehat{C}_{b}\right)+\frac{\left|\widehat{C}_{a_{2}}\right|}{\left|\widehat{C}_{a_{1}}\right|+\left|\widehat{C}_{a_{2}}\right|}\rho\left(\widehat{C}_{a_{2}},\widehat{C}_{b}\right)\ =\ (1+o(1))p_{a_{1}b}.$
So $\rho\left(\widehat{C}_{a_{1}\cup a_{2}},\widehat{C}_{b}\right)=(1+o(1))p_{a_{1}b}$, and repeating this argument, it follows by induction that the average-linkage procedure correctly recovers the tree. ∎
Let us now state and prove the following auxiliary lemma.
Lemma 2.
Consider an HSBM with the same assumptions as in Theorem 1, and let $\widehat{\mathcal{C}}$ be an almost exact estimator of $\mathcal{C}$ ($\widehat{\mathcal{C}}$ is possibly correlated with the graph edges). Then, for any $a,b\in\mathcal{L}_{\mathcal{T}}$ we have $\rho\left(\widehat{C}_{a},\widehat{C}_{b}\right)=(1+o(1))p_{ab}$.
Proof.
We denote by $Z\in\{0,1\}^{N\times K}$ (resp., by $\hat{Z}$) the one-hot representation of the true communities $\mathcal{C}$ (resp., of the predicted communities $\widehat{\mathcal{C}}$), that is $Z_{ic}=\mathbb{1}(i\in C_{c})$ and $\hat{Z}_{ic}=\mathbb{1}(i\in\widehat{C}_{c})$ for all $i\in[N]$ and $c\in\mathcal{L}_{\mathcal{T}}$ (where $\mathbb{1}(\cdot)$ denotes the indicator function). Let $a\neq b\in\mathcal{L}_{\mathcal{T}}$. We shorten the edge density $\rho\left(\widehat{C}_{a},\widehat{C}_{b}\right)$ by $\hat{p}_{ab}$. From the definition of the edge density in (2.1), we have
$\hat{p}_{ab}\ =\ \frac{w\left(\widehat{C}_{a},\widehat{C}_{b}\right)}{\left|\widehat{C}_{a}\right|\cdot\left|\widehat{C}_{b}\right|}\ =\ \frac{\sum_{i\in\widehat{C}_{a},j\in\widehat{C}_{b}}A_{ij}}{\left|\widehat{C}_{a}\right|\cdot\left|\widehat{C}_{b}\right|}\ =\ \frac{\sum_{i,j}\hat{Z}_{ia}\hat{Z}_{jb}A_{ij}}{\sum_{i,j}\hat{Z}_{ia}\hat{Z}_{jb}}.$
Therefore, a variance-bias decomposition leads to
$\left|\hat{p}_{ab}-p_{ab}\right|$ $\ \leq\ \underbrace{\left|\frac{\sum_{i,j}\hat{Z}_{ia}\hat{Z}_{jb}\left(A_{ij}-\mathbb{E}A_{ij}\right)}{\sum_{i,j}\hat{Z}_{ia}\hat{Z}_{jb}}\right|}_{E_{1}}+\underbrace{\left|\frac{\sum_{i,j}\hat{Z}_{ia}\hat{Z}_{jb}\mathbb{E}A_{ij}}{\sum_{i,j}\hat{Z}_{ia}\hat{Z}_{jb}}-p_{ab}\right|}_{E_{2}}.$ (A.1)
Moreover, because $\widehat{\mathcal{C}}$ is almost exact, we have $\sum_{i,j}\hat{Z}_{ia}\hat{Z}_{jb}=(1+o(1))|C_{a}|\cdot|C_{b}|$. Thus,
$\sum_{i,j}\hat{Z}_{ia}\hat{Z}_{jb}\ =\ (1+o(1))\pi_{a}\pi_{b}N^{2},$
using the concentration of multinomial random variables.
Let us bound the two terms $E_{1}$ and $E_{2}$ on the right-hand side of (A.1) separately. To handle the first term, we will use [17, Lemma 4.1]. Let $\bar{p}=2N^{-1}(N-1)^{-1}\sum_{i<j}\operatorname{Var}(A_{ij})$. Because $\operatorname{Var}(A_{ij})=\Theta(\delta_{N})$, we have $\bar{p}=\Theta(\delta_{N})$. Moreover, because $\delta_{N}=\omega(N^{-1})$, we have $\bar{p}\geq 9N^{-1}$ for $N$ large enough. Therefore [17, Lemma 4.1] ensures that with probability at least $1-e^{3}5^{-N}$ we have
[…]
Applying Grothendieck’s inequality [17, Theorem 3.1], we obtain
[…]
where $c$ is Grothendieck’s constant, verifying $20<c<$ 2. Using (A.2) with $X_{i}=\hat{Z}_{ia}$ and $Y_{j}=\hat{Z}_{jb}$ leads to
$E_{1}\ =\ O\left(\sqrt{\frac{\bar{p}}{N}}\right)\ =\ o(\delta_{N}),$
because $p=\Theta(\delta_{N})$ and $N^{-1}=o(\delta_{N})$.
To handle the second term $E_{2}$ in the right-hand side of (A.1), we first notice that
$E_{2}$ $\ =\ \frac{\left|\sum\limits_{i,j}\hat{Z}_{ia}\hat{Z}_{jb}\left(\mathbb{E}A_{ij}-p_{ab}\right)\right|}{\sum\limits_{i,j}\hat{Z}_{ia}\hat{Z}_{jb}}.$ (A.3)
Moreover, for all $i,j\in[N]$ we have
$\mathbb{E}A_{ij}=\sum_{a^{\prime},b^{\prime}\in\mathcal{L}_{\mathcal{T}}}p_{a^{\prime}b^{\prime}}Z_{ia^{\prime}}Z_{jb^{\prime}},$
because $\mathbb{E}A_{ij}=p_{a^{\prime}b^{\prime}}$ if $i\in C_{a^{\prime}}$ and $j\in C_{b^{\prime}}$. Thus,
$\mathbb{E}A_{ij}-p_{ab}\ =\ \sum_{\begin{subarray}{c}a^{\prime},b^{\prime}\in\mathcal{L}_{\mathcal{T}}\\ (a^{\prime},b^{\prime})\neq(a,b)\end{subarray}}(p_{a^{\prime}b^{\prime}}-p_{ab})Z_{ia^{\prime}}Z_{jb^{\prime}}.$
Let $c=\max_{a^{\prime},b^{\prime}}\frac{p_{a^{\prime}b^{\prime}}}{p_{ab}}$. Under our assumptions, $c=\Theta(1)$. We bound the numerator appearing in Equation (A.3) as follows:
$\left|\sum\limits_{i<j}\hat{Z}_{ia}\hat{Z}_{jb}\left(\mathbb{E}A_{ij}-p_{ab}\right)\right|$ $\ \leq\ p_{ab}\left(c+1\right)\sum_{\begin{subarray}{c}a^{\prime},b^{\prime}\in\mathcal{L}_{\mathcal{T}}\\ (a^{\prime},b^{\prime})\neq(a,b)\end{subarray}}\sum_{i<j}\hat{Z}_{ia}\hat{Z}_{jb}Z_{ia^{\prime}}Z_{jb^{\prime}}$
$\ \leq\ 2cp_{ab}\sum_{\begin{subarray}{c}a^{\prime},b^{\prime}\in\mathcal{L}_{\mathcal{T}}\\ (a^{\prime},b^{\prime})\neq(a,b)\end{subarray}}\left|\widehat{C}_{a}\cap C_{a^{\prime}}\right|\cdot\left|\widehat{C}_{b}\cap C_{b^{\prime}}\right|,$
since $c\geq 1$. Let us denote $V\backslash C_{a}=\cup_{a^{\prime}\neq a}C_{a^{\prime}}$ by $C_{a}^{c}$. We have
$\sum_{\begin{subarray}{c}a^{\prime},b^{\prime}\in\mathcal{L}_{\mathcal{T}}\\ (a^{\prime},b^{\prime})\neq(a,b)\end{subarray}}\left|\widehat{C}_{a}\cap C_{a^{\prime}}\right|\cdot\left|\widehat{C}_{b}\cap C_{b^{\prime}}\right|$ $\ =\ \sum_{a^{\prime}\neq a}\sum_{b^{\prime}\in\mathcal{L}_{\mathcal{T}}}\left|\widehat{C}_{a}\cap C_{a^{\prime}}\right|\cdot\left|\widehat{C}_{b}\cap C_{b^{\prime}}\right|+\left|\widehat{C}_{a}\cap C_{a}\right|\sum_{b^{\prime}\neq b}\left|\widehat{C}_{b}\cap C_{b^{\prime}}\right|$
$\ =\ \left|\widehat{C}_{a}\cap C^{c}_{a}\right|\cdot\left|\widehat{C}_{b}\right|+\left|\widehat{C}_{a}\cap C_{a}\right|\cdot\left|\widehat{C}_{b}\cap C_{b}^{c}\right|$
$\ \leq\ \left(\left|\widehat{C}_{b}\right|+\left|\widehat{C}_{a}\right|\right)\mathrm{loss}\left(\mathcal{C},\widehat{\mathcal{C}}\right),$
where the last line uses $|\widehat{C}_{a}\cap C_{a}|\leq|\widehat{C}_{a}|$ and $\mathrm{loss}(\mathcal{C},\widehat{\mathcal{C}})=\sum_{c=1}^{K}\left|\widehat{C}_{c}\cap C^{c}_{c}\right|+\left|\widehat{C}_{c}^{c}\cap C_{c}\right|$. Hence,
[…]
The assumption that $\widehat{\mathcal{C}}$ is almost exact ensures that $\mathrm{loss}(\mathcal{C},\widehat{\mathcal{C}})=o(N)$ and $|\widehat{C}_{a}|,|\widehat{C}_{b}|=\Theta(N)$. Therefore, $E_{2}=o(p_{ab})$. This concludes the proof. ∎
A.2 Proof of Section 3.3
In this section, we prove the Theorem 2 and Proposition 1. We first start in Section A.2.1 by stating a lemma that gives a general condition for the robustness of average-linkage with respect to errors present in $\widetilde{\mathcal{C}}$. We then show in Sections A.2.2 and A.2.3 how to prove Theorem 2 and Proposition 1 from this Lemma 3. The main ingredients of the proof of Lemma 3 are given in Section A.2.4, whereas the most tedious computations are detailed in Section A.2.5.
A.2.1 A general lemma
Let us denote by $B(h)$ the number of bottom communities that have similarity $h$ with a bottom community $a$. By symmetry of BTSBMs, this value $B(h)$ does not depend on $a$, and we have
$B(h)$ =$\ =\ \begin{cases}2^{d-h-1}&h\in\{0,1,\cdots,d-1\},\\ 1&h=d.\end{cases}$ (A.4)
We notice for $0\leq h_{1}\leq d$, $B(h)$ satisfies
$\sum_{h=h_{1}}^{d}B(h)\ =\ B(h_{1}-1),$ (A.5)
where by a slight abuse of notation we define $B(-1)=2^{d}$. The value of $B(h)$ is also equal to the number of bottom communities whose similarity with the bottom community $a$ is no less than $h+1$. Denote by $\bar{p}_{h}$ the expected edge density inside the super community whose similarity with the bottom community $a$ is no less than $h$. By symmetry of BTSBMs, the value of $\bar{p}_{h}$ does not depend on $a$, and we have
$\bar{p}_{h}\ =\ \frac{\sum_{s=h}^{d}B(s)p(s)}{B(h-1)}.$ (A.6)
Let us now state the following lemma.
Lemma 3.
Let $G$ be a BTSBM whose latent binary tree $\mathcal{T}$ has depth $d_{\mathcal{T}}\geq 2$. Suppose that $\pi_{a}=1/K$ for all $a\in\mathcal{L}_{\mathcal{T}}$ and that $\min\limits_{u\in\mathcal{T}}p(u)=\omega(N^{-2})$. Let $\widetilde{\mathcal{C}}$ be a clustering obtained from $\mathcal{C}$ as described in Section 3.3, and $\widehat{\mathcal{T}}$ be the hierarchical tree obtained by average-linkage from $\widetilde{\mathcal{C}}$. Then $\widehat{\mathcal{T}}=\mathcal{T}$ iff for all $h_{ac}\leq d-2$ the following condition is satisfied
$\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\left(\zeta(h_{1})-\zeta(h_{ac})\right)\left(\zeta(h_{2})-\zeta(h_{ac})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$+$ $\ \sum_{h_{1}=h_{ac}+1}^{d-1}\left(B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)\right)+\left(p_{d-1}-p_{h_{ac}}\right)\left(\zeta(d)-\zeta(h_{ac})\right)^{2}$
$+$ $\quad 2\left(p_{d}-p_{d-1}\right)\left(\zeta(d-1)-\zeta(h_{ac})\right)\left(\zeta(d)-\frac{\zeta(d-1)+\zeta(h_{ac})}{2}\right)$
$>$ $\ 0$ (A.7)
The proof of Lemma 3 is tedious, and we defer it to Section A.2.4. Let us first establish Theorem 2 and Proposition 1 from Lemma 3.
A.2.2 Proof of Theorem 2
Proof of Theorem 2 as a corollary of Lemma 3.
Because of assortativity, $p_{h_{2}}-p_{h_{1}}>0$ for all $h_{2}>h_{1}$ and $\bar{p}_{h}>p_{h_{ac}}$ for all $h>h_{ac}$. Moreover, for all $h_{1}<h_{2}$ such that $\zeta(h_{1})\ \leq\ \zeta(h_{2})$, we have $\zeta(h)-\zeta(h_{ac})\geq 0$. Therefore, for all $h_{ac}\geq d-2$, we have
$\sum_{h_{1}=h_{ac}+1}^{d-1}$ $\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\left(\zeta(h_{1})-\zeta(h_{ac})\right)\left(\zeta(h_{2})-\zeta(h_{ac})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
[…]
because each term in the left-hand side of the inequality is non-negative.
Moreover, because $\zeta(d)>\frac{\zeta(d-1)+\zeta(h_{ac})}{2}$,
$\left(p_{d}-p_{d-1}\right)\left(\zeta(d-1)-\zeta(h_{ac})\right)\left(\zeta(d)-\frac{\zeta(d-1)+\zeta(h_{ac})}{2}\right)>0,$
and thus Condition (3) is satisfied. ∎
A.2.3 Proof of Proposition 1
Proof of Proposition 1 as a corollary of Lemma 3.
When $h_{ac}>0$, by substituting $\zeta(h)$ with Equation (3.2), the left hand side of Condition (3) can be written as
$\left(p_{d-1}-p_{h_{ac}}\right)\left(1-\eta\right)^{2}>0,$
which therefore satisfies Condition (3).
When $h_{ac}=0$, we can replace $\zeta(h)$ with Equation (3.2) and denote $\frac{\eta}{2^{d-1}}$ by $\zeta$ for convenience. This allows us to express the left-hand side of Condition (3) as follows:
$\underbrace{\sum_{h_{1}=1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d-1}2B(h_{1})B(h_{2})\zeta^{2}\left(p_{h_{1}}-p_{0}\right)}_{A_{1}}-\underbrace{\sum_{h_{1}=1}^{d-1}2B(h_{1})\zeta\left(1-\eta-\zeta\right)\left(p_{h_{1}}-p_{0}\right)}_{A_{2}}$
$+$ $\ \underbrace{\sum_{h_{1}=1}^{d-1}B^{2}(h_{1})\zeta^{2}\left(\bar{p}_{h_{1}+1}-p_{0}\right)}_{A_{3}}+\left(p_{d-1}-p_{0}\right)\left(1-\eta-\zeta\right)^{2}-2\left(p_{d}-p_{d-1}\right)\zeta\left(1-\eta-\frac{\zeta}{2}\right).$ (A.8)
We notice that
$A_{1}$ $\ =\ \sum_{h_{1}=1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\zeta^{2}\left(p_{h_{1}}-p_{0}\right)-\sum_{h_{1}=1}^{d-1}2B(h_{1})B(d)\zeta^{2}\left(p_{h_{1}}-p_{0}\right),$
$=^{(a)}\sum_{h_{1}=1}^{d-1}2B^{2}(h_{1})\zeta^{2}\left(p_{h_{1}}-p_{0}\right)-\sum_{h_{1}=1}^{d-1}2B(h_{1})B(d)\zeta^{2}\left(p_{h_{1}}-p_{0}\right),$ (A.9)
where in (a), we use Equation (A.5). We also have
$A_{3}$ $=\sum_{h_{1}=1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}B(h_{1})B(h_{2})\zeta^{2}\left(p_{h_{2}}-p_{0}\right),$
$=^{(a)}\sum_{h_{2}=1}^{d}\sum_{h_{1}=1}^{h_{2}-1}B(h_{1})B(h_{2})\left(p_{h_{2}}-p_{0}\right)\zeta^{2},$
$=\sum_{h_{2}=1}^{d}\left(\sum_{h_{1}=1}^{d}B(h_{1})-\sum_{h_{1}=h_{2}}^{d}B(h_{1})\right)B(h_{2})\left(p_{h_{2}}-p_{0}\right)\zeta^{2},$
$=^{(b)}\sum_{h_{2}=1}^{d}\left(B(0)-B(h_{2}-1)\right)B(h_{2})\left(p_{h_{2}}-p_{0}\right)\zeta^{2},$
$=^{(c)}\sum_{h_{1}=1}^{d}B(0)B(h_{1})\zeta^{2}\left(p_{h_{1}}-p_{0}\right)-\left(\sum_{h_{1}=1}^{d-1}2B^{2}(h_{1})\zeta^{2}\left(p_{h_{1}}-p_{0}\right)+\zeta^{2}\left(p_{d}-p_{0}\right)\right)$ (A.10)
where in (a), we swap sums from $\sum_{h_{1}=1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}$ to $\sum_{h_{2}=1}^{d}\sum_{h_{1}=1}^{h_{2}-1}$, in (b), we use Equation (A.5). Finally, to obtain the last equality (c), we re-index $h_{2}$ to $h_{1}$ and use
$B(h-1)\ =\ \begin{cases}2B(h)&h\in\{0,1,\cdots,d-1\},\\ B(d)&h=d,\end{cases}$ (A.11)
(Recall also that $B(d)=1$.) By combining Equations (A.9) and (A.10), we obtain
$A_{1}+A_{2}+A_{3}$ $\ =\ \sum_{h_{1}=1}^{d-1}2B^{2}(h_{1})\zeta^{2}\left(p_{h_{1}}-p_{0}\right)-\sum_{h_{1}=1}^{d-1}2B(h_{1})\zeta\left(1-\eta\right)\left(p_{h_{1}}-p_{0}\right)$
$\qquad+\sum_{h_{1}=1}^{d}B(0)B(h_{1})\zeta^{2}\left(p_{h_{1}}-p_{0}\right)-\left(\sum_{h_{1}=1}^{d-1}2B^{2}(h_{1})\zeta^{2}\left(p_{h_{1}}-p_{0}\right)+\zeta^{2}\left(p_{d}-p_{0}\right)\right),$
$=-\sum_{h_{1}=1}^{d-1}2B(h_{1})\zeta\left(1-\eta\right)\left(p_{h_{1}}-p_{0}\right)+\sum_{h_{1}=1}^{d}B(0)B(h_{1})\zeta^{2}\left(p_{h_{1}}-p_{0}\right)-\zeta^{2}\left(p_{d}-p_{0}\right),$
$=-\sum_{h_{1}=1}^{d}2B(h_{1})\zeta\left(1-\eta\right)\left(p_{h_{1}}-p_{0}\right)+2\zeta\left(1-\eta\right)(p_{d}-p_{0})$
$\qquad+\sum_{h_{1}=1}^{d}B(0)B(h_{1})\zeta^{2}\left(p_{h_{1}}-p_{0}\right)-\zeta^{2}\left(p_{d}-p_{0}\right),$
$=^{(a)}-2B(0)\zeta\left(1-\eta\right)\left(\bar{p}_{1}-p_{0}\right)+2\zeta\left(1-\eta\right)(p_{d}-p_{0})+B^{2}(0)\zeta^{2}\left(\bar{p}_{1}-p_{0}\right)-\zeta^{2}\left(p_{d}-p_{0}\right),$
$=-\eta\left(1-\eta\right)\left(\bar{p}_{1}-p_{0}\right)+2\zeta\left(1-\eta\right)(p_{d}-p_{0})+\eta^{2}\left(\bar{p}_{1}-p_{0}\right)-\zeta^{2}\left(p_{d}-p_{0}\right),$
where in (a) we use Equations (A.6) and (A.11), and in the last equality we use $B(0)\zeta=\eta$.
By substituting $A_{1}+A_{2}+A_{3}$, we can rewrite Equation (A.2.3) as
$\left(\eta^{2}-2\eta\left(1-\eta\right)\right)\left(\bar{p}_{1}-p_{0}\right)+2\zeta\left(1-\eta\right)(p_{d}-p_{0})-\zeta^{2}\left(p_{d}-p_{0}\right)$
$\qquad+\left(p_{d-1}-p_{0}\right)\left(\left(1-\eta\right)^{2}+\zeta^{2}-2\zeta\left(1-\eta\right)\right)-\left(p_{d}-p_{d-1}\right)\zeta\left(2(1-\eta)-\zeta\right),$
$\ =\ \left(\eta^{2}-2\eta\left(1-\eta\right)\right)\left(\bar{p}_{1}-p_{0}\right)+\left(p_{d-1}-p_{0}\right)\left(1-\eta\right)^{2}.$
Therefore, Condition (3) is equivalent to
$\left(\eta^{2}-2\eta\left(1-\eta\right)\right)\left(\bar{p}_{1}-p_{0}\right)+\left(p_{d-1}-p_{0}\right)\left(1-\eta\right)^{2}>0.$ (A.12)
If $p_{d-1}>\bar{p}_{1}$ we have
$\left(\eta^{2}-2\eta\left(1-\eta\right)\right)\left(\bar{p}_{1}-p_{0}\right)+\left(p_{d-1}-p_{0}\right)\left(1-\eta\right)^{2}$ $\ =\ (\bar{p}_{1}-p_{0})\left((1-\eta)^{2}+\eta^{2}-2\eta(1-\eta)\right),$
$\ =\ (\bar{p}_{1}-p_{0})\left(1-2\eta\right)^{2},$
and this quantity is positive since $\bar{p}_{1}>p_{0}$. Therefore, $p_{d-1}>\bar{p}_{1}$ is a sufficient condition to recover the tree, and this proves point (1) of Proposition 1.
If $p_{d-1}\leq\bar{p}_{1}$ we can rewrite the left-hand side of Equation (A.12) (by “completing the square”) as
$(p_{d-1}+3\bar{p}_{1}-4p_{0})\eta^{2}-2(p_{d-1}+\bar{p}_{1}-2p_{0})\eta+(p_{d-1}-p_{0}),$
$\ =\ (p_{d-1}+3\bar{p}_{1}-4p_{0})\left(\eta-\frac{p_{d-1}+\bar{p}_{1}-2p_{0}}{p_{d-1}+3\bar{p}_{1}-4p_{0}}\right)^{2}+(p_{d-1}-p_{0})-\frac{(p_{d-1}+\bar{p}_{1}-2p_{0})^{2}}{p_{d-1}+3\bar{p}_{1}-4p_{0}},$
$\ =\ (p_{d-1}+3\bar{p}_{1}-4p_{0})\left(\eta-\frac{p_{d-1}+\bar{p}_{1}-2p_{0}}{p_{d-1}+3\bar{p}_{1}-4p_{0}}\right)^{2}-\frac{(\bar{p}_{1}-p_{d-1})(\bar{p}_{1}-p_{0})}{p_{d-1}+3\bar{p}_{1}-4p_{0}}.$
Therefore, Equation (A.12) is equivalent to
$(p_{d-1}+3\bar{p}_{1}-4p_{0})\left(\eta-\frac{p_{d-1}+\bar{p}_{1}-2p_{0}}{p_{d-1}+3\bar{p}_{1}-4p_{0}}\right)^{2}-\frac{(\bar{p}_{1}-p_{d-1})(\bar{p}_{1}-p_{0})}{p_{d-1}+3\bar{p}_{1}-4p_{0}}>0.$
Because $(p_{d-1}+3\bar{p}_{1}-4p_{0})>0$, this inequality is satisfied if $\eta<\eta_{-}$ or if $\eta>\eta_{+}$, where
$\eta_{-}$ $\ =\ \frac{p_{d-1}+\bar{p}_{1}-2p_{0}-\sqrt{(\bar{p}_{1}-p_{d-1})(\bar{p}_{1}-p_{0})}}{p_{d-1}+3\bar{p}_{1}-4p_{0}},$
$\eta_{+}$ $\ =\ \frac{p_{d-1}+\bar{p}_{1}-2p_{0}+\sqrt{(\bar{p}_{1}-p_{d-1})(\bar{p}_{1}-p_{0})}}{p_{d-1}+3\bar{p}_{1}-4p_{0}}.$
Notice that $\eta_{+}=\frac{\sqrt{\bar{p}_{1}-p_{d-1}}(2\sqrt{\bar{p}_{1}-p_{0}}-\sqrt{\bar{p}_{1}-p_{d-1}})}{2(p_{d-1}+3\bar{p}_{1}-4p_{0})}+\frac{1}{2}>\frac{1}{2}$, and hence the condition $\eta>\eta_{+}$ cannot be verified (recall that $20\leq\eta<1/$ 2). In contrast, we have $\frac{p_{d-1}+\bar{p}_{1}-2p_{0}}{p_{d-1}+3\bar{p}_{1}-4p_{0}}-\frac{1}{2}=\frac{p_{d-1}-\bar{p}_{1}}{2(p_{d-1}+3\bar{p}_{1}-4p_{0})}\leq 0$ and hence $\eta_{-}\leq\frac{1}{2}$. Moreover, using
$(p_{d-1}+\bar{p}_{1}-2p_{0})^{2}-(\bar{p}_{1}-p_{d-1})(\bar{p}_{1}-p_{0})$ $\ =\ p_{d-1}^{2}+4p_{0}^{2}+3\bar{p}_{1}p_{d-1}-5p_{d-1}p_{0}-3p_{0}\bar{p}_{1}$
= $\ =\ 3(p_{d-1}-p_{0})(\bar{p}_{1}-p_{0})+(p_{d-1}-p_{0})^{2},$
$p_{d-1}+\bar{p}_{1}-2p_{0}>0$, and $(\bar{p}_{1}-p_{d-1})(\bar{p}_{1}-p_{0})>0$, it follows that $\eta_{-}>0$. Therefore, when $p_{d-1}\leq\bar{p}_{1}$, condition (3) is satisfied if $0\leq\eta<\eta_{-}$. ∎
A.2.4 Proof of Lemma 3
Proof of Lemma 3.
We proceed by showing that the edge density $\rho\left(\tilde{C}_{a},\tilde{C}_{b}\right)$ is concentrated around its mean. We denote by $O_{ab}=C_{a}\cap\tilde{C}_{b}$ the nodes in cluster $C_{a}$ but assigned to cluster $\tilde{C}_{b}$. In particular, the set of all misclassified nodes $O$ is given by $O=\cup_{\begin{subarray}{c}a,b\in\mathcal{L}_{\mathcal{T}}\\ b\neq a\end{subarray}}O_{ab}$ and is independent of the edges. Therefore, $w\left(O_{ab},O_{cd}\right)\sim\operatorname{Bin}\left(|O_{ab}|\cdot|O_{cd}|,p_{ac}\right)$. We note, from the definition of the edge density (2.1) and the fact that $\tilde{C}_{a}=\cup_{k\in\mathcal{L}_{\mathcal{T}}}O_{ka}$, that
$\rho\left(\tilde{C}_{a},\tilde{C}_{b}\right)\ =\ \frac{w\left(\tilde{C}_{a},\tilde{C}_{b}\right)}{\left|\tilde{C}_{a}\right|\cdot\left|\tilde{C}_{b}\right|}\quad\text{ with }\quad w\left(\tilde{C}_{a},\tilde{C}_{b}\right)\ =\ \sum_{k,\ell\in\mathcal{L}_{\mathcal{T}}}w\left(O_{ka},O_{\ell b}\right),$ (A.13)
We will express $\mathbb{E}w\left(O_{ka},O_{\ell b}\right)$ for each scenario. The concentration $\rho\left(\tilde{C}_{a},\tilde{C}_{b}\right)$ around its mean then follows from Chernoff’s bound.
In the remainder of the proof, we let $a,b,c\in\mathcal{L}_{\mathcal{T}}$ be three different leaves such that $b$ is the closest leaf from $a$, i.e., $b=\operatorname*{arg\,max}\limits_{k\in\mathcal{L}_{\mathcal{T}}}|\mathrm{lca}(a,k)|$ and $c\not\in\{a,b\}$. To conclude that the linkage procedure outputs the correct tree, we have to verify that
$\forall c\not\in\{a,b\}\colon\rho\left(\tilde{C}_{a},\tilde{C}_{b}\right)>\rho\left(\tilde{C}_{a},\tilde{C}_{c}\right).$
Since $\rho\left(\tilde{C}_{a},\tilde{C}_{c}\right)=(1+o(1))\rho\left(\tilde{C}_{b},\tilde{C}_{c}\right)$, Equation (2.2) ensures $\rho\left(\tilde{C}_{a\cup b},\tilde{C}_{c}\right)=(1+o(1))\rho\left(\tilde{C}_{a},\tilde{C}_{c}\right)$ after the merge, which in turn guarantees that the new super community $a\cup b$ does not merge with any another community before all the bottom communities are merged. After all the bottom communities are merged, we again obtain balanced (super) communities and repeat these steps for these super communities as we did for the bottom communities.
The random variables $w\left(O_{ka},O_{\ell b}\right)$ follow a Binomial distribution such that $\mathbb{E}w\left(O_{ka},O_{\ell b}\right)\ =\ \left(\frac{N}{K}\right)^{2}\zeta(|\mathrm{lca}(k,a)|)\zeta(|\mathrm{lca}(\ell,b)|)p_{|\mathrm{lca}(k,\ell)|}$, where $\zeta(h)$ is the fraction of nodes mislabeled from a bottom community with similarity $h$ to bottom community $a$. We define $h_{1}=|\mathrm{lca}(k,a)|$, $h_{2}=|\mathrm{lca}(\ell,b)|$, and $h_{ac}=|\mathrm{lca}(a,c)|$.
When $h_{1}=h_{ac}<h_{2}\leq d$, each node of the fraction $\zeta(h_{1})=\zeta(h_{ac})$ of the bottom cluster $c$ is mislabeled as a node inside a uniformly randomly chosen bottom cluster whose similarity with the bottom cluster $a$ is no less than $h_{1}+1=h_{ac}+1$. Furthermore, each node of the fraction $\zeta(h_{2})$ of the bottom cluster $a$ is mislabeled as a node inside a uniformly randomly chosen bottom cluster whose similarity with the bottom cluster $a$ is $h_{2}~{}(<h_{1}=h_{ac})$. Therefore, in this case, the similarity between a node from $a$ and a node from $c$ is exactly the same as when the two nodes are randomly picked from super community $C_{u}:~{}C_{u}\ =\ \bigcup_{b\in\mathcal{L}_{\mathcal{T}}\colon b_{1:(h_{ac}+1)}=a_{1:(h_{ac}+1)}}C_{b}$. Hence, the expected edge probability between these two nodes is $\bar{p}_{h_{1}+1}=\bar{p}_{h_{ac}+1}$. Similarly, when $h_{1}=h_{2}<h_{ac}$, each node of the fraction $\zeta(h_{1})=\zeta(h_{h_{2}})$ of both bottom clusters $a$ and $c$ is mislabeled as a node inside a uniformly randomly chosen bottom cluster whose similarity with the bottom cluster $a$ (and also with $c$) is $h_{1}=h_{2}$. Therefore, in this case, the expected edge probability between a node from $a$ and a node from $c$ is $\bar{p}_{h_{1}+1}=\bar{p}_{h_{2}+1}$. In the other cases, the similarity between bottom clusters $a$ and $c$ is $\min(p_{h_{1}},p_{h_{2}},p_{h_{ac}})=\min(p_{h_{1}},p_{h_{ac}})$.
Now we obtain whp that for any $c\neq a$,
$\mathbb{E}w\left(O_{ka},O_{\ell c}\right)\ =\ \left(\frac{N}{K}\right)^{2}\times\begin{cases}\zeta(h_{1})\zeta(h_{2})p_{h_{1}}&\text{ if }0\leq h_{1}<h_{ac}\text{ and }h_{1}<h_{2}\leq d,\\ \zeta(h_{ac})\zeta(h_{2})\bar{p}_{h_{ac}+1}&\text{ if }h_{1}=h_{ac}\text{ and }h_{ac}<h_{2}\leq d,\\ \zeta(h_{1})\zeta(h_{2})p_{h_{ac}}&\text{ if }h_{ac}<h_{1}<h_{2}\leq d,\\ \zeta^{2}(h_{1})\bar{p}_{h_{1}+1}&\text{ if }0\leq h_{1}=h_{2}<h_{ac},\\ \zeta^{2}(h_{1})p_{h_{ac}}&\text{ if }h_{ac}\leq h_{1}=h_{2}\leq d,\\ \end{cases}$
The special case $h_{ac}=d-1$ occurs when $c=b$. In that case, we have
$\mathbb{E}w\left(O_{ka},O_{\ell b}\right)\ =\ \left(\frac{N}{K}\right)^{2}\times\begin{cases}\zeta(h_{1})\zeta(h_{2})p_{h_{1}}&\text{ if }0\leq h_{1}<d-1\text{ and }h_{1}<h_{2}\leq d,\\ \zeta(d-1)\zeta(d)p_{d}&\text{ if }(h_{1},h_{2})=(d-1,d),\\ \zeta^{2}(h_{1})\bar{p}_{h_{1}+1}&\text{ if }h_{1}=h_{2}<d-1,\\ \zeta^{2}(h_{1})p_{d-1}&\text{ if }d-1\leq h_{1}=h_{2}\leq d,\\ \end{cases}$
Because $c\notin\{a,b\}$, we have $h_{ac}<d-1$. Thus, $\mathbb{E}w\left(O_{ka},O_{\ell b}\right)-\mathbb{E}w\left(O_{ka},O_{\ell c}\right)$ is equal to
$\left(\frac{N}{K}\right)^{2}\times\begin{cases}0&\text{ if }0\leq h_{1}<h_{ac}\text{ and }h_{1}<h_{2}\leq d,\\ \zeta(h_{ac})\zeta(h_{2})\left(p_{h_{ac}}-\bar{p}_{h_{ac}+1}\right)&\text{ if }h_{1}=h_{ac}<h_{2}\leq d,\\ \zeta(h_{1})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)&\text{ if }h_{ac}<h_{1}<h_{2}\leq d\text{ and }h_{1}<d-1,\\ \zeta(d-1)\zeta(d)\left(p_{d}-p_{h_{ac}}\right)&\text{ if }(h_{1},h_{2})=(d-1,d),\\ 0&\text{ if }0\leq h_{1}=h_{2}<h_{ac},\\ \zeta^{2}(h_{1})\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)&\text{ if }h_{ac}\leq h_{1}=h_{2}<d-1,\\ \zeta^{2}(h_{1})\left(p_{d-1}-p_{h_{ac}}\right)&\text{ if }d-1\leq h_{1}=h_{2}\leq d.\\ \end{cases}$
Because $B(h)$ is the number of bottom communities having a similarity $h$ with the bottom community $a$, we obtain
$\rho\left(\tilde{C}_{a},\tilde{C}_{b}\right)-\rho\left(\tilde{C}_{a},\tilde{C}_{c}\right)$
[…]
$\qquad\qquad\quad+\sum_{h_{1}=h_{ac}+1}^{d-2}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{1})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)+2\zeta(d-1)\zeta(d)\left(p_{d}-p_{h_{ac}}\right)$
$\qquad\qquad\quad+\sum_{h_{1}=h_{ac}}^{d-1}B^{2}(h_{1})\zeta^{2}(h_{1})\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)+\zeta^{2}(d)\left(p_{d-1}-p_{h_{ac}}\right)\Bigg{)}.$ (A.14)
We show in Section A.2.5 that we can further express $\rho\left(\tilde{C}_{a},\tilde{C}_{b}\right)-\rho\left(\tilde{C}_{a},\tilde{C}_{c}\right)$ as
$(1+o(1))\Bigg{(}\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\left(\zeta(h_{1})-\zeta(h_{ac})\right)\left(\zeta(h_{2})-\zeta(h_{ac})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\qquad\qquad\qquad\qquad+\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)$
$\qquad\qquad\qquad\qquad+2\left(\zeta(d-1)-\zeta(h_{ac})\right)\left(\zeta(d)-\frac{\zeta(d-1)+\zeta(h_{ac})}{2}\right)\left(p_{d}-p_{d-1}\right)$
$\qquad\qquad\qquad\qquad+\left(\zeta(d)-\zeta(h_{ac})\right)^{2}\left(p_{d-1}-p_{h_{ac}}\right)\Bigg{)}.$ (A.15)
Therefore, the condition $\rho\left(\tilde{C}_{a},\tilde{C}_{b}\right)-\rho\left(\tilde{C}_{a},\tilde{C}_{c}\right)>0$ is equivalent to condition (3). ∎
A.2.5 Additional computation for the proof of Lemma 3
In this subsection, we detail the tedious computations that allow us to transform Equation (A.2.4) into Equation (A.2.4). for the proof of Lemma 3. In order to highlight the differences between the lines, we sometimes use bold characters. Let $\Delta\rho=\rho\left(\tilde{C}_{a},\tilde{C}_{b}\right)-\rho\left(\tilde{C}_{a},\tilde{C}_{c}\right)$. From Equation (A.2.4), we have
$\Delta\rho\ =$ $(1+o(1))\Bigg{(}-\sum_{h_{2}=h_{ac}+1}^{d}2B(h_{ac})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(\bar{p}_{h_{ac}+1}-p_{h_{ac}}\right)$
$\qquad\qquad\quad+\ \sum_{h_{1}=h_{ac}+1}^{d-2}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{1})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\qquad\qquad\quad+2B(d-1)B(d)\zeta(d-1)\zeta(d)\left(p_{d}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}}^{d-2}B^{2}(h_{1})\zeta^{2}(h_{1})\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)$
$\qquad\qquad\quad+\sum_{h_{1}=d-1}^{d}B^{2}(h_{1})\zeta^{2}(h_{1})\left(p_{d-1}-p_{h_{ac}}\right)\Bigg{)}$
[…]
$\qquad\qquad\quad+\sum_{h_{1}=h_{ac}+1}^{d-2}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{1})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\qquad\qquad\quad+2\zeta(d-1)\zeta(d)\left(p_{d}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}}^{d-2}\sum_{h=h_{1}+1}^{d}B(h_{1})\boldsymbol{B(h)}\zeta^{2}(h_{1})\left(\boldsymbol{p_{h}}-p_{h_{ac}}\right)$
$\qquad\qquad\quad+\left(\zeta^{2}(d-1)+\zeta^{2}(d)\right)\left(p_{d-1}-p_{h_{ac}}\right)\Bigg{)}.$
Thus,
$\Delta\rho\ =$ $(1+o(1))\Bigg{(}-\sum_{h_{1}=h_{ac}+1}^{d}\sum_{h_{2}=h_{ac}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\qquad\qquad\quad+\sum_{h_{1}=h_{ac}+1}^{\boldsymbol{d-1}}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{1})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\qquad\qquad\quad\boldsymbol{-\ 2\zeta(d-1)\zeta(d)\left(p_{d-1}-p_{h_{ac}}\right)}+2\zeta(d-1)\zeta(d)\left(p_{d}-p_{h_{ac}}\right)$
$\qquad\qquad\quad+\sum_{h_{1}=h_{ac}}^{\boldsymbol{d-1}}\sum_{h=h_{1}+1}^{d}B(h_{1})B(h)\zeta^{2}(h_{1})\left(p_{h}-p_{h_{ac}}\right)\boldsymbol{-\zeta^{2}(d-1)\left(p_{d}-p_{h_{ac}}\right)}$
$\qquad\qquad\quad+\left(\zeta^{2}(d-1)+\zeta^{2}(d)\right)\left(p_{d-1}-p_{h_{ac}}\right)\Bigg{)},$
[…]
$\qquad\qquad\quad+\underbrace{\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{1})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)}_{T_{2}}$
$\qquad\qquad\quad+\underbrace{\sum_{h_{1}=h_{ac}}^{d-1}\sum_{h=h_{1}+1}^{d}B(h_{1})B(h)\zeta^{2}(h_{1})\left(p_{h}-p_{h_{ac}}\right)}_{T_{3}}+\boldsymbol{\zeta^{2}(d)\left(p_{d-1}-p_{h_{ac}}\right)}$
$\qquad\qquad\quad+\boldsymbol{\left(2\zeta(d-1)\zeta(d)-\zeta^{2}(d-1)\right)\left(p_{d}-p_{d-1}\right)}\Bigg{)}.$ (A.16)
We compute $T_{1}$ and $T_{3}$ separately. We have
$T_{1}$ $\ =\ \sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d}2B^{2}(h_{1})\zeta(h_{ac})\zeta(h_{1})\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d}\sum_{h_{2}=h_{ac}+1}^{h_{1}-1}2B(h_{1})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right),$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d}2B^{2}(h_{1})\zeta(h_{ac})\zeta(h_{1})\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\boldsymbol{\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}}2B(h_{1})\zeta(h_{ac})\zeta(h_{1})\left(\boldsymbol{B(h_{2})}p_{h_{2}}-\boldsymbol{B(h_{2})}p_{h_{ac}}\right),$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d}2B^{2}(h_{1})\zeta(h_{ac})\zeta(h_{1})\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d-1}2\boldsymbol{B^{2}(h_{1})}\zeta(h_{ac})\zeta(h_{1})\left(\boldsymbol{\bar{p}_{h_{1}+1}}-p_{h_{ac}}\right),$
and
$T_{3}$ $\ =\ \sum_{h_{1}=h_{ac}+1}^{d}\sum_{h_{2}=h_{ac}}^{h_{1}-1}B(h_{2})B(h_{1})\zeta^{2}(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right),$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d}\sum_{h_{2}=\boldsymbol{h_{ac}+1}}^{h_{1}-1}B(h_{1})B(h_{2})\zeta^{2}(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)\boldsymbol{+\sum_{h_{1}=h_{ac}+1}^{d}B(h_{ac})B(h_{1})\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right)},$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d}\sum_{h_{2}=h_{ac}+1}^{h_{1}-1}B(h_{1})B(h_{2})\zeta^{2}(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d}\boldsymbol{\sum_{h_{2}=h_{ac}+1}^{d}}B(h_{1})\boldsymbol{B(h_{2})}\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right),$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d}\sum_{h_{2}=h_{ac}+1}^{h_{1}-1}B(h_{1})B(h_{2})\zeta^{2}(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d}\sum_{h_{2}=h_{ac}+1}^{\boldsymbol{h_{1}-1}}B(h_{1})B(h_{2})\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d}\boldsymbol{B^{2}(h_{1})}\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{\boldsymbol{h_{2}=h_{1}+1}}^{d}B(h_{1})B(h_{2})\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right).$
We can further manipulate $T_{3}$ to obtain
$T_{3}$ $\ =\ \sum_{h_{1}=h_{ac}+1}^{d}\sum_{h_{2}=h_{ac}+1}^{h_{1}-1}B(h_{1})B(h_{2})\left(\zeta^{2}(h_{2})+\zeta^{2}(h_{ac})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d}B^{2}(h_{1})\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}B(h_{1})B(h_{2})\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right),$
$\ =\ \boldsymbol{\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}}B(h_{1})\left(\zeta^{2}(h_{1})+\zeta^{2}(h_{ac})\right)\left(\boldsymbol{B(h_{2})}p_{h_{2}}-\boldsymbol{B(h_{2})}p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d}B^{2}(h_{1})\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}B(h_{1})B(h_{2})\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right),$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d-1}\boldsymbol{B^{2}(h_{1})}\left(\zeta^{2}(h_{1})+\zeta^{2}(h_{ac})\right)\left(\boldsymbol{\bar{p}_{h_{1}+1}}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d}B^{2}(h_{1})\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}B(h_{1})B(h_{2})\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right).$
Combining the expressions for $T_{1}$ and $T_{3}$, we obtain
$-T_{1}+T_{3}$
$\ =\ -\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)-\sum_{h_{1}=h_{ac}+1}^{d}2B^{2}(h_{1})\zeta(h_{ac})\zeta(h_{1})\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad-\sum_{h_{1}=h_{ac}+1}^{d-1}2B^{2}(h_{1})\zeta(h_{ac})\zeta(h_{1})\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta^{2}(h_{1})+\zeta^{2}(h_{ac})\right)\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d}B^{2}(h_{1})\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}B(h_{1})B(h_{2})\zeta^{2}(h_{ac})\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}B(h_{1})B(h_{2})\left(\zeta^{2}(h_{ac})-2\zeta(h_{ac})\zeta(h_{2})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d}B^{2}(h_{1})\left(\zeta^{2}(h_{ac})-2\zeta(h_{ac})\zeta(h_{1})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right).$
Furthermore,
$-T_{1}+T_{3}$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}B(h_{1})B(h_{2})\left(\zeta^{2}(h_{ac})-2\zeta(h_{ac})\zeta(h_{2})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{\boldsymbol{d-1}}\boldsymbol{\sum_{h_{2}=h_{1}+1}^{d}}B(h_{1})\boldsymbol{B(h_{2})}\left(\zeta^{2}(h_{ac})-2\zeta(h_{ac})\zeta(h_{1})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\boldsymbol{B^{2}(d)\left(\zeta^{2}(h_{ac})-2\zeta(h_{ac})\zeta(d)\right)\left(p_{d}-p_{h_{ac}}\right)}+\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right),$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}B(h_{1})B(h_{2})\left(2\zeta^{2}(h_{ac})-2\zeta(h_{ac})\zeta(h_{2})-2\zeta(h_{ac})\zeta(h_{1})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\left(\zeta^{2}(h_{ac})-2\zeta(h_{ac})\zeta(d)\right)\left(p_{d}-p_{h_{ac}}\right)+\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right),$
and
$-T_{1}+T_{3}-\left(p_{d}-p_{h_{ac}}\right)\left(\zeta^{2}(h_{ac})-2\zeta(h_{ac})\zeta(d)\right)+T_{2}$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}B(h_{1})B(h_{2})\left(2\zeta^{2}(h_{ac})-2\zeta(h_{ac})\zeta(h_{2})-2\zeta(h_{ac})\zeta(h_{1})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{1})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right),$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\left(\zeta(h_{1})\zeta(h_{2})+\zeta^{2}(h_{ac})-\zeta(h_{ac})\zeta(h_{2})-\zeta(h_{ac})\zeta(h_{1})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right),$
$\ =\ \sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\boldsymbol{\left(\zeta(h_{1})-\zeta(h_{ac})\right)\left(\zeta(h_{2})-\zeta(h_{ac})\right)}\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\quad+\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right).$
Finally we come back to Equation (A.16) and obtain
$\Delta\rho$ $\ =\ (1+o(1))\Big{(}-T_{1}+T_{2}+T_{3}+\zeta^{2}(d)\left(p_{d-1}-p_{h_{ac}}\right)$
$\qquad\qquad\qquad\quad+\left(2\zeta(d-1)\zeta(d)-\zeta^{2}(d-1)\right)\left(p_{d}-p_{d-1}\right)\Big{)},$
$\ =\ (1+o(1))\Big{(}\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\left(\zeta(h_{1})-\zeta(h_{ac})\right)\left(\zeta(h_{2})-\zeta(h_{ac})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\qquad\qquad\qquad\quad+\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)+\zeta^{2}(d)\left(p_{d-1}-p_{h_{ac}}\right)$
$\qquad\qquad\qquad\quad+\left(\zeta^{2}(h_{ac})-2\zeta(h_{ac})\zeta(d)\right)\left(p_{d}-p_{h_{ac}}\right)+\left(2\zeta(d-1)\zeta(d)-\zeta^{2}(d-1)\right)\left(p_{d}-p_{d-1}\right)\Big{)},$
$\ =\ (1+o(1))\Big{(}\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\left(\zeta(h_{1})-\zeta(h_{ac})\right)\left(\zeta(h_{2})-\zeta(h_{ac})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\qquad\qquad\qquad\quad+\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)+\zeta^{2}(d)\left(p_{d-1}-p_{h_{ac}}\right)$
$\qquad\qquad\qquad\quad+\left(\zeta^{2}(h_{ac})-2\zeta(h_{ac})\zeta(d)\right)\left(\left(p_{d}\boldsymbol{-p_{d-1}}\right)+\left(\boldsymbol{p_{d-1}}-p_{h_{ac}}\right)\right)$
$\qquad\qquad\qquad\quad+\left(2\zeta(d-1)\zeta(d)-\zeta^{2}(d-1)\right)\left(p_{d}-p_{d-1}\right)\Big{)},$
$=(1+o(1))\Bigg{(}\sum_{h_{1}=h_{ac}+1}^{d-1}\sum_{h_{2}=h_{1}+1}^{d}2B(h_{1})B(h_{2})\left(\zeta(h_{1})-\zeta(h_{ac})\right)\left(\zeta(h_{2})-\zeta(h_{ac})\right)\left(p_{h_{1}}-p_{h_{ac}}\right)$
$\qquad\qquad\qquad\quad+\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)$
$\qquad\qquad\qquad\quad+2\left(\zeta(d-1)-\zeta(h_{ac})\right)\left(\zeta(d)-\frac{\zeta(d-1)+\zeta(h_{ac})}{2}\right)\left(p_{d}-p_{d-1}\right)$
$\qquad\qquad\qquad\quad+\left(\zeta(d)-\zeta(h_{ac})\right)^{2}\left(p_{d-1}-p_{h_{ac}}\right)\Bigg{)},$
and this last expression establishes Equation (A.2.4).
Appendix B Proofs of Sections 4 and 5
B.1 Proof of Theorem 3
Let us first recall some notations and results for the general SBM. Let $\pi\in(0,1)^{K}$ be a probability vector and $p\in(0,1)^{K\times K}$ a symmetric matrix. We denote $G\sim{\sf SBM}(N,\pi,p)$ if
  • each node $i\in[N]$ of $G$ is assigned to a unique community $C_{k}$, with $k\sim\mathrm{Multi}(1,[K],\pi)$;
  • two nodes $i\in C_{k}$ and $j\in C_{\ell}$ are connected with probability $p_{k\ell}$.
In the following, we denote by $\mathcal{A}=\{A_{1},\cdots,A_{t}\}$ a collection of $t$ non-empty and non-overlapping subsets of $[K]$ such that $\cup_{r=1}^{t}A_{r}=[K]$. An algorithm exactly recovers $\mathcal{A}$ if it assigns each node $i$ in $G$ to an element of $\{A_{1},\cdots,A_{t}\}$ that contains its true community (up to a relabelling of the $A_{r}$’s) with probability $1-o(1)$. For two non-overlapping subsets $A,B\subset[K]$, we denote by $\mathrm{CH}\left(A,B\right)=\mathrm{CH}(A,B,\pi,p)$ the quantity
$\mathrm{CH}\left(A,B\right)\ =\ \min_{\begin{subarray}{c}a\in A\\ b\in B\end{subarray}}\,\sup_{t\in(0,1)}(1-t)\sum_{c=1}^{K}\pi_{c}\operatorname{D}_{t}\left(\operatorname{Ber}(p_{ab})\|\operatorname{Ber}(p_{bc})\right),$
and by $I(\mathcal{A})=I(\mathcal{A},\pi,p)$ the quantity
$I(\mathcal{A})\ =\ \min_{A_{r}\neq A_{t}\in\mathcal{A}}\mathrm{CH}\left(A_{r},A_{t}\right).$
We define the finest partition of $[K]$ with threshold $\tau$ the partition $\mathcal{A}^{*}_{\tau}$ of $[K]$ such that
$\mathcal{A}^{*}_{\tau}\ =\ \operatorname*{arg\,max}_{\mathcal{A}}\left\{\left|\mathcal{A}\right|\colon I(\mathcal{A})>\tau\right\}.$
It is the partition of $[K]$ in the largest number of subsets, among all partitions that verify $I(\mathcal{A})>\tau$. The following theorem holds.
Theorem 4.
Let $G\sim{\sf SBM}(N,\pi,p)$ and $\mathcal{A}$ a partition of $[K]$ in $t$ non-empty and non-overlapping elements. Suppose that no two rows of $p$ are equal. Note that if $\tau$ is too large, such a partition might not exist. The following holds.
  1. (i)
    if $I(\mathcal{A})<(1-\Omega(1))\,\,N^{-1}\log N$, then no algorithm can exactly recover $\mathcal{A}$;
  2. (ii)
    the agnostic-degree-profiling algorithm of [3] exactly recovers the finest partition with threshold $\tau=(1+\Omega(1))\,\,N^{-1}\log N$.
Proof.
The proof of point (i) can be found in [2, Theorem 1] while point (ii) corresponds to [3, Theorem 4]. ∎
Finally, we have the following lemma for the estimation of the link probabilities $p_{ab}$.
Lemma 4.
Let $\widehat{\mathcal{C}}$ be an exact estimator of $\mathcal{C}$ and suppose that $\min_{u\in\mathcal{T}}p(u)=\omega(N^{-2})$. Then with high probability $\rho(\widehat{C}_{a},\widehat{C}_{b})=(1+o(1))p_{ab}$.
Proof.
We assume that the permutation $\tau$ in the definition of the loss function (Equation (3.1)) is the identity. Furthermore, we shorten $\rho(\widehat{C}_{a},\widehat{C}_{b})$ by $\hat{p}_{ab}$. Since $\widehat{\mathcal{C}}$ is an exact estimator of $\mathcal{C}$, we have for $N$ large enough that $C_{a}=\widehat{C}_{a}$ for all $a\in\mathcal{L}_{\mathcal{T}}$. Hence,
$\hat{p}_{ab}\ =\ \frac{\sum_{i\in C_{a},j\in C_{b}}A_{ij}}{|C_{a}|\cdot|C_{b}|}.$
Since $\sum_{i\in C_{a},j\in C_{b}}A_{ij}\sim\operatorname{Bin}(|C_{a}|\cdot|C_{b}|,p_{ab})$ and $p_{ab}=\omega(N^{-2})$ as well as $|C_{a}|,|C_{b}|=\Theta(N)$, we conclude that $\hat{p}_{ab}=(1+o(1))p_{ab}$ using the concentration of binomial distribution. ∎
We can now prove Theorem 3.
Proof of Theorem 3.
Since HSBM is a special instance of the general SBM, in which the communities are indexed by elements of $\mathcal{L}_{\mathcal{T}}$ instead of elements of $[K]$, we can directly apply Theorem 4. The set $\mathcal{S}_{q}$ of nodes at depth $q$ naturally forms a partition $\mathcal{A}=(A_{t})_{t\in\mathcal{S}_{q}}$ of $\mathcal{L}_{\mathcal{T}}$ as follows: $\ell\in A_{t}$ iff $\mathrm{lca}(t,\ell)=t$. Exactly recovering this partition is equivalent to recovering exactly $\mathrm{sc}(q,\mathcal{C},\mathcal{T})$, and therefore point (i) of Theorem 3 follows from point (i) of Theorem 4.
Conversely, point (ii) of Theorem 4 implies that we can recover the finest partition $q^{*}$. In the case of an HSBM, this corresponds to the largest $q$ such that $I_{q}>(1+\Omega(1))\,N^{-1}\log N$. Since $q\mapsto I_{q}$ is non-decreasing we have $q\geq q^{*}$. Moreover, using Lemma 4, we proceed as in the proof of Theorem 1 to show that we recover the tree $\mathcal{T}[\mathcal{S}_{\leq q^{*}}]$. In particular, we can recover the super-communities at any higher depth $q\geq q^{*}$, which are exactly the depths verifying $I_{q}>(1+\Omega(1))\,N^{-1}\log N$. ∎
B.2 Proof of Lemma 1
Proof of Lemma 1.
Let $a,b$ be two leaves of $\mathcal{T}$ such that their least common ancestor is at a depth $s$ strictly less than $q$, i.e., $|\mathrm{lca}(a,b)|=s<q$. For any leaf $c\in\mathcal{L}_{T}$ we have $p_{ac}=p_{bc}$ if $c\not\in\mathcal{L}_{\mathcal{T}[u]}$. Therefore, the sum in (4.1) can be limited to $c\in\mathcal{L}_{\mathcal{T}[u]}$ so that
$\mathrm{CH}\left(a,b\right)$ $\ =\ \frac{1}{K}\sup_{t\in(0,1)}(1-t)\sum_{c\in\mathcal{L}_{\mathcal{T}[u]}}\operatorname{D}_{t}\left(\operatorname{Ber}(p_{ac})\|\operatorname{Ber}(p_{bc})\right).$
In the following, we let $u=\mathrm{lca}(a,b)$ and $s=|u|$. For any two nodes $v,w\in\mathcal{T}$, we denote by $\mathrm{sim}(v,w)$ the depth of the least common ancestor to $v$ and $w$, that is $\mathrm{sim}(v,w)=|\mathrm{lca}(v,w)|$. Finally, we denote by $\Gamma_{a,b}(v,w)$ the set of leaves of $\mathcal{T}[u]$ for which the common ancestor with $a$ is $v$ and the common ancestor with $b$ is $w$, i.e.,
$\Gamma_{a,b}(v,w)\ =\ \left\{c\in\mathcal{L}_{\mathcal{T}[u]}\colon\mathrm{lca}(a,c)=v\text{ and }\mathrm{lca}(b,c)=w\right\}.$
We have
$\left|\Gamma_{a,b}(v,w)\right|\ =\ \begin{cases}1&\text{ if }\mathrm{sim}(a,v)=d\text{ and }\mathrm{sim}(b,w)=s\text{ (this is equivalent to $v=a$ and $w=u$)},\\ 1&\text{ if }\mathrm{sim}(a,v)=s\text{ and }\mathrm{sim}(b,w)=d\text{ (this isequivalent to $v=u$ and $w=b$)},\\ 2^{k-1}&\text{ if }\mathrm{sim}(a,v)=d-k\text{ and }\mathrm{sim}(b,w)=s\text{ for some }k\in[d-s-1],\\ 2^{k-1}&\text{ if }\mathrm{sim}(a,v)=s\text{ and }\mathrm{sim}(b,w)=d-k\text{ for some }k\in[d-s-1],\\ 0&\text{ otherwise.}\end{cases}$
Moreover, for any $c\in\Gamma_{a,b}(v,w)$ we have $p_{ac}=p(\mathrm{lca}(a,c))=p_{\mathrm{sim}(a,v)}$ and similarly $p_{bc}=p_{\mathrm{sim}(b,w)}$. Thus,
$\mathrm{CH}(a,b)\ =\ \frac{1}{K}\sup_{t\in(0,1)}(1-t)\left\{\operatorname{D}_{t}\left(p_{d}\|p_{s}\right)+\operatorname{D}_{t}\left(p_{s}\|p_{d}\right)+\sum_{k=1}^{d-s-1}2^{k-1}\left[\operatorname{D}_{t}\left(p_{d-k}\|p_{s}\right)+\operatorname{D}_{t}\left(p_{s}\|p_{d-k}\right)\right]\right\},$ (B.1)
where we shortened $\operatorname{D}_{t}(\operatorname{Ber}(p)\|\operatorname{Ber}(q))$ by $\operatorname{D}_{t}(p,q)$.
Finally, let $P,Q$ be two probability distributions. By the concavity of $t\mapsto(1-t)\operatorname{D}_{t}(P\|Q)$ and the relation $(1-t)\operatorname{D}_{t}(P\|Q)=t\operatorname{D}_{1-t}(Q\|P)$ (see for example [34]), the function inside the $\sup$ of Equation B.1 is concave and symmetric around $t=1/2$. Therefore the $\sup$ of Equation B.1 is achieved at $t=1/2$. Hence,
$\mathrm{CH}(a,b)\ =\ \frac{1}{K}\left[\operatorname{D}_{1/2}\left(p_{s},p_{d}\right)+\sum_{k=1}^{d-s-1}2^{k-1}\operatorname{D}_{1/2}\left(p_{s},p_{d-k}\right)\right]\ =\ H_{s}.$
This last quantity $H_{s}$ depends on $a,b$ only via their similarity $\mathrm{sim}(a,b)=s$. Moreover, since the network is assortative we have $p_{s}<p_{s+1}$ and thus $\operatorname{D}_{1/2}\left(p_{s},p_{\ell}\right)>\operatorname{D}_{1/2}\left(p_{s+1},p_{\ell}\right)$ for $\ell\geq s+1$. Therefore $s\mapsto H_{s}$ is decreasing, and consequently
$I_{q}\ =\ \min_{\begin{subarray}{c}a\neq b\in\mathcal{L}_{\mathcal{T}}\\ \mathrm{sim}(a,b)<q\end{subarray}}\mathrm{CH}(a,b)\ =\ \min_{s<q}H_{s}\ =\ H_{q-1}.$
B.3 Comparing top-down versus bottom-up conditions
Lemma 5.
For any $q\in\{1,\cdots,d-1\}$ we have $J_{q}^{\mathrm{bu}}>J_{q}^{\mathrm{td}}$. When $q=d$, we have $J_{d}^{\mathrm{bu}}=J_{d}^{\mathrm{td}}$.
Proof of Lemma 5.
We have
$2^{d}J_{q}^{\mathrm{bu}}$ $\ =\ a_{q}+2^{d-q}a_{q-1}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}-2\sqrt{a_{q-1}}\left(\sqrt{a_{d}}+\sum_{k=1}^{d-q}2^{k-1}\sqrt{a_{d-k}}\right),$
$2^{d}J_{q}^{\mathrm{td}}$ $\ =\ a_{q}+2^{d-q}a_{q-1}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}-2\sqrt{2^{d-q}a_{q-1}}\sqrt{a_{d}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}}.$
Hence,
$2^{d}\left(J_{q}^{\mathrm{bu}}-J_{q}^{\mathrm{td}}\right)$ = $\ =\ 2\sqrt{a_{q-1}}\left(2^{\frac{d-q}{2}}\sqrt{a_{d}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}}-\left(\sqrt{a_{d}}+\sum_{k=1}^{d-q}2^{k-1}\sqrt{a_{d-k}}\right)\right)$
= $\ =\ 2\sqrt{a_{q-1}}(D-E)$
$\ =\ \frac{2\sqrt{a_{q-1}}}{D+E}(D^{2}-E^{2}),$
where
$D\ =\ 2^{\frac{d-q}{2}}\sqrt{\left(a_{d}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}\right)}\quad\text{ and }\quad E\ =\ \sqrt{a_{d}}+\sum_{k=1}^{d-q}2^{k-1}\sqrt{a_{d-k}}.$
Since $\frac{2\sqrt{a_{q-1}}}{D+E}>0$, we focus on showing $D\geq E$.
First, we notice that when $d=q$, $D=E=\sqrt{a_{d}}$ and hence $J_{d}^{\mathrm{td}}=J_{d}^{\mathrm{bu}}$.
Next, when $d-q\geq 1$, we have
$D^{2}-E^{2}$ = $\ =\ 2^{d-q}\left(a_{d}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}\right)-\left(\sqrt{a_{d}}+\sum_{k=1}^{d-q}2^{k-1}\sqrt{a_{d-k}}\right)^{2}$
$\ =\ (2^{d-q}-1)a_{d}+\sum_{k=1}^{d-q}2^{k-1}\left(2^{d-q}-2^{k-1}\right)a_{d-k}-\sum_{k=1}^{d-q}2^{k}\sqrt{a_{d}}\sqrt{a_{d-k}}-\sum_{k=1}^{d-q-1}\sum_{l>k}^{d-q}2^{k+l-1}\sqrt{a_{d-k}}\sqrt{a_{d-l}}.$
Noticing that $\sum_{k=1}^{d-q}2^{k-1}=2^{d-q}-1$ and that $\sum_{l=1,l\neq k}^{d-q}2^{k+l-2}=2^{k-1}\left(\sum_{l=1}^{d-q}2^{l-1}-2^{k-1}\right)=2^{k-1}(2^{d-q}-2^{k-1}-1)$ leads to
$D^{2}-E^{2}\ =$ $\left(\sum_{k=1}^{d-q}2^{k-1}a_{d}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}-2\sum_{k=1}^{d-q}2^{k-1}\sqrt{a_{d}}\sqrt{a_{d-k}}\right)$
[…]
Using the network’s assortativity, we conclude that this last quantity is strictly greater than zero, and hence $J_{q}^{\mathrm{bu}}>J_{q}^{\mathrm{td}}$ for all $q\leq d-1$. ∎
Appendix C Additional numerical experiments
C.1 Synthetic data sets
C.1.1 Unbalanced HSBM
We also examine the accuracy of HCD algorithms on HSBM whose binary tree is not necessarily full and balanced. Similarly to the BTSBM, we still suppose that the depth of the tree $\mathcal{T}$ determines the link probabilities, i.e., $p(u)=p_{|u|}$ for all $u\in\mathcal{T}$. Moreover, we assume that the primitive communities have the same number of nodes. Figures 9 amd 10 compare the different HCD algorithms on HSBM whose respective unbalanced trees are given in Figures 9(a) and 10(a). We observe that bottom-up and top-down perform similarly well, but one sometimes outperforms the other depending on $\beta$ and the metric used. The relative error for the estimation of the tree similarity matrix is defined as $\|S(\widehat{\mathcal{T}},\widehat{\mathcal{C}})-S(\mathcal{T},\mathcal{C})\|^{2}_{F}/\|S(\mathcal{T},\mathcal{C})\|^{2}_{F}$, where the tree-similarity matrix $S\left(\mathcal{T},\mathcal{C}\right)$ is defined in (6.1).
(a) Unbalanced tree
(b) AMI
(c) Tree similarity
(d) Estimated $K$
Figure 9: (a) Hierarchical tree. (b)-(d) Results of the different HCD algorithms on unbalanced HSBMs whose hierarchical tree is given in Figure 9(a), and with $N=3200$ nodes and $p_{k}=64\cdot\beta^{5-k}\cdot\frac{\log{N}}{N}$. The results are averaged over 100 realizations, and error bars (vertical solid lines) show the standard error but are typically smaller than the symbols.
(a) Unbalanced tree
(b) AMI
(c) Tree similarity
(d) Estimated $K$
Figure 10: (a) Hierarchical tree. (b)-(d) Results of the different HCD algorithms on unbalanced BTSBMs whose tree shape is given in Figure 10(a) and with $N=3200$ nodes and $p_{k}=144\cdot\beta^{5-k}\cdot\frac{\log{N}}{N}$. The results are averaged over 100 realizations, and error bars (vertical solid lines) show the standard error but are typically smaller than the symbols.
C.2 Real data sets
C.2.1 Military inter-alliance
We next consider the network of military alliances between countries. The data is provided by the Alliance Treaty Obligations and Provisions (ATOP) project [22]. We select the year 2018 (as this is the most recent year available). We define two countries as allied if they share a defensive alliance [^12] In particular, we do not consider non-aggression pacts, as those are more numerous and historically not necessarily well respected.. This leads to a network of 133 countries and 1391 alliances. Some important countries such as India or Switzerland are missing as they do not share any defensive alliances with anybody. Moreover, the graph is not connected, as a small component made of three countries (China, North Korea and Cuba) is disconnected from the rest of the world.
Figures 11 and 12 show the output of bottom-up and top-down algorithms. Bottom-up predicts 7 bottom communities, which represent geopolitical alliances based on political affiliation and geography (European countries, Eurasian countries, Arabic countries, Western African countries and Central/Southern African countries). The top level splits the graph’s largest connected component into 3 clusters: Western countries, Eurasian countries, and African and Middle-East countries. While some of these clusters are also recovered by top-down HCD, the separation of African countries by top-down algorithm appears worse.
(a) Highest depth.
(b) Middle depth (after 2 merges).
(c) Lowest depth (after 3 merges).
(d) Dendrogram.
Figure 11: Output of bottom-up algorithm on the military alliance network. The dendrogram does not show the disconnected component (China, Cuba, North Korea).
(a) Highest depth (8 clusters).
(b) Middle depth (6 clusters).
(c) Lowest depth (3 clusters).
(d) Dendrogram.
Figure 12: Output of top-down algorithm on the military alliance network. The dendrogram does not show the disconnected component (China, Cuba, North Korea).
C.2.2 Football data set
We also test HCD algorithms on the United States college (American) football dataset [16]. This network represents the schedule of Division I games for the 2000 season. Each node in the network corresponds to a college football team, and edges represent the regular-season games between the teams. The teams are categorized into 11 conferences, in which games are more frequent between the members. Each conference has 8 to 12 teams. We exclude the ”independent” teams which do not belong to any conferences. Since the original community labels appear to be based on the 2001 season, while the edges represent the games played during the 2000 season, we proceed to the same correction as in [13].
The results obtained by the different HCD algorithms are given in Figure 13. First, we observe that all the algorithms perform well (AMI for bottom-up, top-down, Paris, and Bayesian are respectively 0.962, $0.892$, 0.965, and 0.976). However, top-down has more errors than the other algorithms. Interestingly, bottom-up, top-down, and Bayesian algorithms predict 10 clusters (more precisely, Bayesian detects 10.1 communities averaged over 100 runs), as they tend to infer Big West and Mountain West conferences as a single cluster. Finally, we can restore some geographical proximity among conferences from the hierarchy inferred by bottom-up. For example, Conference USA is composed of teams located in the Southern United States, while the Southeastern Conference’s member institutions are located primarily in the South Central and Southeastern United States. Another example is that teams belonging to Pacific Ten, Big West, and Mountain West are all located in the West, and these conferences are also close in the bottom-up dendrogram.
(a) bottom-up
(b) top-down
Figure 13: Output of bottom-up and top-down algorithms on the football data set. The colours correspond to conferences, and grey edges indicate having regular-season games between the two teams. The hierarchical tree is drawn in black, and its root is marked by a star symbol.